### v3

# LightGaussian: Unbounded 3D Gaussian Compression with 15x Reduction and 200+ FPS

Abstract
Recent advancements in real-time neural rendering using point-based techniques have paved the way for the widespread adoption of 3D representations. However, foundational approaches like 3D Gaussian Splatting come with a substantial storage overhead caused by growing the SfM points to millions, often demanding gigabyte-level disk space for a single unbounded scene, posing significant scalability challenges and hindering the splatting efficiency. To address this challenge, we introduce LightGaussian, a novel method designed to transform 3D Gaussians into a more efficient and compact format. Drawing inspiration from the concept of Network Pruning, LightGaussian identifies Gaussians that are insignificant in contributing to the scene reconstruction and adopts a pruning and recovery process, effectively reducing redundancy in Gaussian counts while preserving visual effects. Additionally, LightGaussian employs distillation and pseudo-view augmentation to distill spherical harmonics to a lower degree, allowing knowledge transfer to more compact representations while maintaining scene appearance. Furthermore, we propose a hybrid scheme, VecTree Quantization, to quantize all attributes, resulting in lower bitwidth representations with minimal accuracy losses. In summary, LightGaussian achieves an averaged compression rate over 15$\times$ while boosting the FPS from 139 to 215, enabling an efficient representation of complex scenes on Mip-NeRF 360, Tank & Temple datasets. Project website: https://lightgaussian.github.io/
Figure 1: Compressibility and Rendering Speed. We present LightGaussian to transform 3D Gaussians into a more compact representation. LightGaussian effectively prunes redundant Gaussians while preserving visual fidelity (on the left). Consequently, it reduces the average storage from 727MB to 42MB and improves the FPS from 139 to 215.
11footnotetext: Equal Contribution.
1 Introduction
Novel view synthesis (NVS) aims to generate photo-realistic images of a 3D scene from unobserved viewpoints, given a set of calibrated multi-view images. NVS holds importance as it can be used for a wide range of real-world applications, including virtual reality [12], augmented reality [73], digital twin [13], and autonomous driving [65]. Neural Radiance Fields (NeRFs)[44, 4, 5] have demonstrated promising ability for photo-realistic 3D modeling and synthesis from multi-view images where 3D location and view directions are mapped to view-dependent color and volumetric density. The pixel intensity can be rendered using the volume rendering technique[14]. However, NeRF and its variants have struggled with rendering speeds for practical deployment in real-world scenarios. Subsequent efforts along this trend introduce voxel-based implicit representation [56, 64, 52, 69, 10], hash grid [46], computational parallelism [49], or neural light field representation [61] to speed up the rendering. However, they either require task-specific design or face a trade-off between rendering quality and efficiency, making them difficult to generalize to large-scale scenarios with practical speed. Recent progress in point-based 3D Gaussian Splatting (3D-GS) [32] has brought photo-realistic rendering quality to the real-time level, even for complex scenes. The authors utilize the idea of representing the scene as explicit 3D Gaussians with specific properties to model the scene; the 2D images are efficiently rendered using a technique named splatting[36]. The optimal balance between speed and quality indicates a potential trend to use 3D-GS as a new representation for generating numerous large-scale scenes in digital twins and autonomous driving. However, using point-based representations inevitably introduces significant storage costs, as each point and its attributes are stored independently. For instance, 3D-GS on a typical unbounded 360 scene [5] requires more than one gigabyte, which inhibits the scalability of 3D-GS (e.g., 1.4GB on scene Bicycle).
In this paper, we counteract the heavy storage issue and deliver a compact representation while preserving the rendering quality. By examining the well-trained point-based representation, each scene is composed of millions of Gaussians, grown from sparse Structure from Motion (SfM) point cloud. Attributes are attached to Gaussians to model the scene’s geometry and appearance. However, the significant number of Gaussians, as well as the high-degree Spherical Harmonics (SH) coefficients used for modeling scene reflection, contribute to an over-parametrized representation when fitting to the scenes. To minimize the required Gaussian number, we propose proper criteria to measure the global significance of each 3D Gaussian, in the context of its contribution for view synthesis. Gaussians with low impact on visual quality will be identified and pruned, followed by short recovery steps to be applied. Spherical harmonics (SH) coefficients, as the majority of the data, are used for modeling the view-dependent color. Compressing it directly by shrinking the higher degree harms the reflectance effect, and we propose a general distillation step enhanced by pseudo-view augmentation to harmlessly transfer the knowledge into a compact level. VecTree Quantization step adaptively picks “just some amount” of distinct point attributes based on the global significance, further reducing the required bitwidth of the original format.
In summary, our proposed framework, LightGaussian, efficiently reduces the Gaussian count (e.g., from 1.49M to 575K, Fig. 1, left), significantly reducing storage requirements from 727MB to 42MB, with minimal rendering quality decrease ($\downarrow$ 0.013 in SSIM) on the Mip-NeRF 360 datasets (Fig. 1, right). LightGaussian further improves the rendering speed to a higher level (200+ FPS) on complex scenes containing detailed backgrounds, suggesting a viable solution for broadening the application scope.
2 Related Works
Figure 2: The overall pipeline of LightGaussian. 3D Gaussians are optimized from multi-view images and SfM points. LightGaussian first calculates the global significance for each Gaussian based on training observations; Gaussians with the least significance are pruned. A subsequent distillation with synthesized pseudo-views is introduced to transfer the SH into a compact format. VecTree quantization, consisting of lossy codebook quantization and lossless octree quantization, is further introduced for compressing the representation.
Efficient 3D Scene Representations for NVS
Neural radiance field (NeRF) [44] uses a multi-layer perceptron (MLP) to represent a scene, and this compact representation has brought view synthesis quality to a new stage. However, NeRF suffers from extremely slow inference challenges. Follow-ups either explore ray re-parameterizations[4, 5], explicit spatial data structures [68, 21, 40, 31, 56, 46, 10], caching and distillation [61, 22, 28, 49], or ray-based representations [2, 55] for speeding up. Still, NeRF-based methods struggle to achieve real-time rendering speed in practical large-scale scenes, which is caused by the multiple queries needed for rendering a single pixel, limiting their practical use. Point-based representations (e.g., Gaussian) have been extensively explored in various domains, including shape reconstruction [33], molecular structures modeling [7], and the replacement of point-cloud data [15]. Additionally, these representations are utilized in shadow [47] and cloud rendering applications [42]. Recently, 3D Gaussian Splatting (3D-GS) [32], combining the idea of point-based rendering and splatting techniques for rendering, achieves real-time speed with comparable rendering quality to the best MLP-based renderer, Mip-NeRF 360 [5]. Although promising, the heavy storage requirement to store all attributes attached to the Gaussians often requires gigabyte-level disk space for saving a single unbounded scene, and millions of Gaussians hinder the rendering efficiency of 3D-GS.
Model Pruning and Vector Quantization
Model pruning involves reducing the complexity of a neural network by eliminating non-significant parameters to maintain a balance between performance and resource utilization. Unstructured [37] and structured pruning [27, 1] remove components at weight-level and neuron (mostly channel) levels to provide a smaller network with a smaller or more efficient network architecture. The iterative magnitude pruning (IMP) method, where weights of the smallest magnitude are progressively pruned over multiple iterations, has been highly successful in lottery ticket rewinding [19, 20]. Additionally, vector quantization [59] aims to represent data with discrete entries of a learned codebook (i.e., tokens) to achieve lossy compression. In general, the mean square error (MSE) is used to find the most similar pattern in the codebook to replace the original input data vector. Previous works [11, 23, 43] have shown that learning a discrete and compact representation not only contributes to visual understanding but also improves the robustness of models. In this vein, vector quantization has been widely adopted in image synthesis [16], text-to-image generation [24], and novel view synthesis [68, 39, 25].
Knowledge Distillation
Knowledge Distillation (KD) is widely adopted in various machine learning tasks [54, 51, 63, 61]. A student model is trained for the primary purpose of model compression [8] or ensembling [30] with the help of teacher models. Many variants [3, 50, 60, 70, 29, 41, 58] have been proposed to achieve better knowledge transfer from the teacher to the student models. In the field of 3D vision, neural scene representations have embraced knowledge distillation mainly through view renderings to leverage existing 2D priors. DreamFusion [48] and NeuralLift-360 [66] adopt pre-trained text-to-image diffusion models for 3D generation. DFF [35], NeRF-SOS [18], INS [17], SA3D [9] distill various pre-trained 2D image feature extractors to perform corresponding tasks in the 3D domain. Knowledge distillation has also played a key role in model compression of scene representations. R2L [61] and KiloNeRF [49] distill Neural Radiance Fields (NeRF) into more efficient representations such as light fields or multiple tiny MLPs. Our work falls into the category of utilizing knowledge distillation for model compression and leverages novel view renderings as the bridge to connect the teacher and student models.
3 Method
Overview
The overview of LightGaussian is illustrated in Fig. 2. The 3D-GS model is trained using multi-view images and is initially initialized from SfM point clouds. By expanding the sparse points to millions of Gaussians, the scene is well-represented. Then, the 3D-GS undergoes processing within our pipeline to transform it into a more compact format. This involves utilizing Gaussian Prune and Recovery to reduce the number of Gaussians, SH Distillation to remove redundant SHs while preserving the modeled specular light, and VecTree Quantization to store Gaussians at a lower bit-width.
3.1 Background: 3D Gaussian Splatting
3D Gaussian Splatting (3D-GS) [32] is an explicit point-based 3D scene representation, utilizing Gaussians with various attributes to model the scene. When representing a complex real-world scene, 3D-GS is initialized from an SfM sparse point cloud, and Gaussian Densifications are applied to increase the Gaussian counts that are used for handling small-scale geometry insufficiently covered. Formally, each Gaussian is characterized by a covariance matrix $\boldsymbol{\Sigma}$ and a center point $\boldsymbol{X}$, which is referred to as the mean value of the Gaussian:
$\text{G}(\boldsymbol{X})=e^{-\frac{1}{2}\boldsymbol{X}^{T}\boldsymbol{\Sigma}^{-1}\boldsymbol{X}},\boldsymbol{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T},$ (1)
where $\boldsymbol{\Sigma}$ can be decomposed into a scaling matrix $\mathbf{S}$ and a rotation matrix $\mathbf{R}$ for differentiable optimization.
The complex directional appearance is modeled by an additional property, Spherical Harmonics (SH), with $n$ coefficients, $\left\{\boldsymbol{c}_{i}\in\mathbb{R}^{3}|i=1,2,\ldots,n\right\}$ where $n=D^{2}$ represents the number of coefficients of SH with degree $D$. A higher degree $D$ equips 3D-GS with a better capacity to model the view-dependent effect but causes a significantly heavier attribute load.
When rendering 2D images from the 3D Gaussians, the technique of splatting [67, 36] is employed for the Gaussians within the camera planes. With a viewing transform denoted as $\boldsymbol{W}$ and the Jacobian of the affine approximation of the projective transformation represented by $\boldsymbol{J}$, the covariance matrix $\boldsymbol{\Sigma}^{\prime}$ in camera coordinates can be computed as follows:
$\boldsymbol{\Sigma}^{\prime}=\boldsymbol{JW}\boldsymbol{\Sigma}\boldsymbol{W}^{T}\boldsymbol{J}^{T}.$ (2)
Specifically, for each pixel, the color and opacity of all the Gaussians are computed using the Gaussian’s representation Eq. 1. The blending of $N$ ordered points that overlap the pixel is given by the formula:
$\boldsymbol{C}=\sum_{i\in N}\boldsymbol{c_{i}}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{i}).$ (3)
Here, $c_{i}$, $\alpha_{i}$ represents the color and opacity of this point computed by a Gaussian with covariance $\boldsymbol{\Sigma}$ multiplied by an optimizable per-point opacity and SH color coefficients. In summary, each Gaussian point is characterized by attributes including: position $\boldsymbol{X}\in\mathbb{R}^{3}$, color defined by spherical harmonics coefficients $\boldsymbol{C}\in\mathbb{R}^{(k+1)^{2}}\times 3$ (where $k$ represents the degrees of freedom), opacity $\alpha\in\mathbb{R}$, rotation factor $\boldsymbol{r}\in\mathbb{R}^{4}$, and scaling factor $\boldsymbol{s}\in\mathbb{R}^{3}$.
3.2 Gaussian Pruning & Recovery
Gaussian densification [32], which involves cloning and splitting the initial SfM point cloud, is employed to address the challenge of insufficient coverage and is used to model small-scale geometry as well as detailed scene appearance. While this strategy leads to significantly improved reconstruction quality, it results in the number of Gaussians growing from thousands to millions after optimization. Such an explicit point-based representation with a large number of Gaussians requires an extremely significant storage overhead. However, pruning the Gaussians based on simplistic criteria (e.g., point opacity) can lead to a substantial degradation in modeling performance, especially where the intricate scene structure may be eliminated, as is demonstrated in Fig. 3.
Figure 3: Zero-shot Opacity-based Pruning. A significant number of Gaussians exhibit small opacity values (top). Simply utilizing Gaussian opacity as an indicator for pruning the least important Gaussians results in the rendered image losing intricate details (bottom), with the PSNR dropping from 27.2 to 25.3. This has inspired us to find better criteria to measure global significance in terms of rendering quality.
Drawing inspiration from the success of representative neural network pruning techniques [26], which eliminate less impactful neurons without compromising the network’s overall performance, we tailor a general pruning paradigm for point-based representation to reduce the over-parameterized point number in a manner that preserves the original accuracy. Therefore, identifying the most representative redundant Gaussians with recoverable accuracy is a crucial step in our approach.
Global Significance Calculation
Simply relying on Gaussian opacity as a significance criterion leads to suboptimal Gaussian pruning, prompting the need for a more effective formula. Inspired by Equation 3, these 3D Gaussians can be rasterized by projecting them onto a specific camera viewpoint for image rendering. The significance of each Gaussian can then be quantified based on its contribution to each pixel across all training views, with similar principle in magnitude network pruning [38]. Consequently, we iterate over all training pixels to calculate the hit count of each Gaussian, factoring in both Gaussian opacity and volume:
$\text{GS}_{\text{j}}$ $=\sum_{i=1}^{MHW}\mathbbm{1}(\text{G}(\boldsymbol{X}_{j}),r_{i})\cdot\sigma_{j}\cdot\gamma_{j}(\boldsymbol{\Sigma}),$ (4)
where $j$ indicates the Gaussian index, and $M$, $H$, and $W$ represent the number of training views, image height, and width, respectively. $\mathbbm{1}$ is the indicator function that determines whether a Gaussian intersects with a given ray.
The use of absolute Gaussian volume tends to exaggerate the importance of background Gaussians, leading to the immoderate pruning of Gaussians that model intricate geometry. Therefore, we propose a more adaptive way to measure the dimension of its volume:
$\gamma(\boldsymbol{\Sigma})$ $=(\text{V}_{\text{norm}})^{\beta},$ (5)
$\text{V}_{\text{norm}}$ $=\text{min}\left(\text{max}\left(\frac{\text{V}(\boldsymbol{\Sigma})}{\text{V}_{\text{max90}}},0\right),1\right).$
Here, the volume is firstly normalized by the 90% largest of all sorted Gaussians, clipping the range between 0 and 1, and $\beta$ is introduced to provide additional flexibility.
Gaussian Co-adaptation
We rank all the Gaussians based on the computed global significance scores, and provide a quantitative basis for pruning Gaussians with the lowest scores. It is worth noting that an overly aggressive pruning ratio can be detrimental to performance. Stripping away too many Gaussians can lead to a noticeable degradation in the model’s accuracy and visual fidelity since incomplete scene coverage. To mitigate this, a phase of short Gaussian co-adaptation is essential, which involves a joint adjustment of the Gaussians’ attributes, allowing them to adapt and compensate for the loss incurred by pruning.
3.3 Distilling into Compact SHs
In the uncompressed Gaussian Splat data, a significant portion is comprised of Spherical Harmonics (SH) coefficients, requiring (45+3) floating-point values per splat, which represent 81.3 percent of the total attribute volume. Reducing the degree of SH, while beneficial for decreasing disk space usage, leads to a noticeable loss of surface ‘shininess’, particularly affecting specular reflection variance when the viewpoint is altered.
To achieve a balance between disk storage efficiency and scene reflectance quality, we propose transferring knowledge from well-trained high-degree SHs to their compact counterparts (lower-degree) via data distillation:
$\mathcal{L}_{\text{distill}}(\text{G}_{\text{SH2-deg}})=\frac{1}{N}\sum_{i=1}^{N}\left\lVert\boldsymbol{C}_{\text{teacher}}-\boldsymbol{C}_{\text{student}}\right\rVert_{2}^{2}.$ (6)
Simply reintroducing these view-dependent visual effects to the network does not significantly enhance its knowledge. We, therefore, suggest employing the concept of data augmentation. This involves training the Gaussians not only to represent known views but also to learn behaviors from unseen (pseudo) views modeled by the teacher model, thus broadening their representational capacity.
Synthesize Pseudo Views
The specular reflection rays are reflected off the surface of the light source when we move the viewpoints. We can augment the training views by sampling additional viewpoints that can reflect such reflectance. As we know the camera locations and viewing directions of all training views, we randomly sample pseudo views around each of the training views, obeying a Gaussian distribution, while fixing the camera view directions:
$\mathbf{t}_{\text{pseudo}}=\mathbf{t}_{\text{train}}+\mathcal{N}(0,\sigma^{2}),$ (7)
where $\mathbf{t}_{\text{pseudo}}$ and $\mathbf{t}_{\text{train}}$ represent the newly synthesized and training camera positions, respectively. $\mathcal{N}$ denotes a Gaussian distribution with mean 0 and variance $\sigma^{2}$, which is added to the original position to generate the new position.
3.4 VecTree Attribute Compression
Vector quantization is well-explored in voxel-based NeRF representations [39, 68, 57, 72]. Although a high compression rate can be reached, Gaussian attributes cannot be easily quantized in the same fashion. The reason is that each attribute of the explicit representation has its own physical meaning, especially when representing large-scale 3D scenes, and attributes such as opacity, rotation, and scale are more sensitive to such discretized representation, with a dramatic drop in accuracy. Therefore, we propose two ways to mitigate this gap: first, we reutilize the significance score from Sec. 3.2 and perform excessive vector quantization on the SHs of the least important Gaussians. For the Gaussian location, we draw from point cloud compression and adopt octree-based lossless compression into our framework. For the remaining important SHs and other attributes representing the Gaussian shape, rotation, and opacity, we save these features in float16 format.
Vector Quantization
The codebook under vector quantization (VQ) is initialized with $K$ codes within it, denoted as $\boldsymbol{C}={\boldsymbol{c}_{1},\boldsymbol{c}_{2},\boldsymbol{c}_{3},\boldsymbol{c}_{K}}$ to represent the Gaussian sets $\boldsymbol{G}={\boldsymbol{g}_{1},\boldsymbol{g}_{2},\boldsymbol{g}_{3},\boldsymbol{g}_{N}}$ where $K\ll N$. To assign each code to the Gaussian feature, we randomly select a batch of Gaussians at each iteration and calculate the Euclidean distance between the selected Gaussian and each code vector in the codebook to determine which code the Gaussian is associated with. Then, a significance-weighted codebook optimization is applied: $\mathbf{c}_{i}:=\lambda\cdot\mathbf{c}_{i}+(1-\lambda)\cdot\sum{\mathbf{g}_{j}\in\mathcal{R}(\mathbf{c}_{k})}\text{GS}_{j}\cdot\mathbf{g}_{j}$, where $\mathbf{g}_{j}\in\mathcal{R}(\mathbf{c}_{i})$ denotes the $j$-th Gaussian assigned to the $i$-th code, and $\lambda$ controls the contribution from the moving average. We only perform excessive VQ on the least 60% of Gaussians, since we are representing complex, large-scale scenes.
Methods Mip-NeRF 360 Datasets Tank and Temple Datasets
[…]
Plenoxels [69] 6.79 2100MB 23.62 0.670 0.443 11.2 2700MB 21.07 0.719 0.379
INGP-Big [46] 9.43 48MB 26.75 0.751 0.299 2.79 48MB 21.92 0.744 0.305
Mip-NeRF 360 [5] 0.06 8.6MB 29.23 0.844 0.207 0.09 8.6MB 22.21 0.759 0.257
VQ-DVGO [39] 6.53 56.07MB 25.44 0.699 0.325 - - - - -
3D-GS [32] 139 727MB 29.13 0.870 0.185 106 380MB 23.11 0.822 0.219
Ours 215 42.48MB 28.45 0.857 0.210 209 22.43MB 22.83 0.807 0.242
Table 1: Quantitative Comparisons in Real-world Large-scale Scenes. We compare LightGaussian with the original 3D-GS [32], efficient voxel-based NeRFs [69, 46], voxel NeRF with vector quantization [39], and compact MLP-based NeRF [5]. Voxel-based methods all exhibit a lack of sufficient capacity for representing large-scale scenes and are not able to run at real-time speed. Mip-NeRF 360 produces the best visual quality, but it requires more than 16s to render a single image. Our method achieves a good balance among FPS, model size, and rendering quality.
Figure 4: Visual Comparisons. We compare LightGaussian with the vanilla 3D-GS [32], presenting a residual map between prediction and ground-truth scaled from 0 to 127 to highlight the differences. We observe that LightGaussian preserves the specular reflections (yellow boxes) after converting to a compact format. Additionally, a slight lightness change is noted after the conversion, as shown in the bottom white box. For dynamic viewpoint comparisons, please refer to our supplementary video material.
Octree-based Compression
Unstructured and high-precision 3D points pose a challenge for efficient storage, especially for large-scale scenes. Applying lossy compression to Gaussian locations shows significant rendering degradation, as the Gaussian location is sensitive to the subsequent rasterization accuracy. Therefore, drawing from the Point Cloud Compression literature [53], which applies lossless encoding on the Gaussian location, utilizing the octree geometry codec in G-PCC [45] consumes only a very small amount of bits. Specifically, let $(X_{i}=(x_{i},y_{i},z_{i}))_{i=1\ldots N}$ be the set of 3D positions associated with the points of the input point cloud. The L-PCC encoder [53] computes the quantized positions $(\hat{X}_{i})_{i=1...N}$ as follows: $\hat{X}_{i}=\left\lfloor(X_{i}-X_{shift})\times s\right\rfloor,$ where $X_{shift}$ and $s$ are pre-defined parameters and we set all octree-related parameters the same as PCGCv2 [62].
For the decoding stage, the reconstructed positions $(\hat{X}_{i})_{i=1}^{N}$ are generated by applying the following inverse quantization process: $\hat{X}_{i}=\frac{\hat{X}i}{s}+X_{shift}$.
4 Experiments
4.1 Experimental Settings
Datasets and Metrics
Comparisons are performed on the widely adopted scene-scale view synthesis data created by Mip-NeRF360 [6], which contains nine real-world large-scale scenes (5 unbounded outdoors, 4 indoors with detailed backgrounds). We utilize the published seven scenes for comparison following the train/test split in Mip-NeRF360 [6]. Additionally, we use another large-scale unbounded dataset, the Tanks and Temples dataset [34], where we perform comparisons using the same two scenes as used in 3D-GS [32]. In total, the adopted datasets consist of scenes with very different capture styles and cover both bounded indoor scenes with detailed backgrounds and large unbounded outdoor cases. The rendering quality is reported using metrics such as peak signal-to-noise ratio (PSNR), structural similarity (SSIM), and perceptual similarity via LPIPS [71].
Compared Baselines
We compare with methods that can model large-scale scenes, including Plenoxel [69], Mip-NeRF360 [5], and 3D-GS [32], along with representative efficient methods such as Instant-ngp [46], which uses a hash grid for storage, and VQ-DVGO [39], utilizing DVGO [56] as the basic representation and introducing voxel pruning and vector quantization to it.
[…]
[1] Baseline (3D-GS [32]) 192.05 353MB 31.68 0.926 0.200
[2]  + Gaussian Pruning. 312.30 116MB 30.32 0.911 0.222
[3]   + Co-adaptation 303.99 116MB 31.85 0.925 0.206
[4]  + SH Compactness. 318.97 77MB 30.54 0.914 0.217
[5]   + Distillation 304.20 77MB 31.47 0.922 0.211
[6]   + Pseudo-views 300.60 77MB 31.59 0.923 0.211
[7]  + Codebook Quant. 300.60 23MB 31.26 0.917 0.220
[8]  + Octree Compression. 300.60 20MB 31.26 0.917 0.220
[9] LightGaussian (Ours) 300.60 20MB 31.26 0.917 0.220
Table 2: Ablation studies on the Gaussian Pruning, SH Compactness, and the VecTree Compression. Scene: Room. Zero-shot Gaussian pruning leads to a degradation in rendering quality (#2), but Co-adaptation can recover most of the scene details (#3). Directly eliminating high-order SH negatively affects the quality (#4), while distillation with pseudo-view helps to mitigate the gap (#5, #6). Codebook quantization significantly reduces the required model size (#7), while lossless Octree compression helps to preserve the overall quality.
[…]
Baseline 192.05 353MB 31.68 0.926 0.200
Hit Count Only 300.57 116MB 28.28 0.895 0.238
$\times$ Opacity. 310.29 116MB 30.10 0.910 0.222
$\times$ Opacity $\times$ $\gamma$(Volume). 312.30 116MB 30.32 0.911 0.222
  + Co-adaptation 303.99 116MB 31.85 0.925 0.206
Table 3: Ablation study of the Gaussian Pruning, by using different Gaussian attributes for computing its global significance score. By considering only the hit count of each Gaussian from training rays, the zero-shot pruning leads to inferior performance. Incorporating the opacity and volume drives us to a better criterion. The subsequent Gaussian Co-adaptation is used to recover most of the information loss from the pruning of redundant Gaussians.
[…]
Baseline 77.09MB 31.59 0.923 0.211
+FP16 36.76MB 31.58 0.923 0.212
 + VQ All att. 18.58MB 22.97 0.750 0.378
 + VQ All att. $\times$ GS 18.58MB 26.39 0.830 0.327
 + VQ SH. 22.53MB 30.94 0.910 0.225
 + VQ SH $\times$ GS 22.53MB 31.26 0.917 0.220
VecTree 20.65MB 31.26 0.917 0.220
Table 4: Ablation study of hybrid VecTree Quantization (VQ). By quantizing all attributes to FP16, except for Gaussian locations due to their sensitivity, a smaller model is achieved. VQ is applied to all attributes leads to inferior modeling accuracy, but this can be mitigated by using Global Significance (GS) on the least crucial Gaussians. Other attributes (e.g., scale) are also sensitive to VQ, hence we only apply VQ on the SH. Combining this with Octree compression results in our VecTree Compression (the last row), which demonstrates a good balance between size and quality.
Implementation Details
The framework is developed using Pytorch and reuses the differential Gaussian rasterization presented by 3D-GS [32]. We use our own trained checkpoints that match performance with those detailed in the 3D-GS paper [32]. For the Global Significance Calculation, we set the power to 0.1 in Eq. 5, and fine-tune for 5,000 steps in Gaussian Co-adaptation. For the SHs Distillation, we reduce the 3-degree SHs to 2-degree, eliminating 21 elements for each Gaussian, and enhance it with $\sigma$ set to 0.1 in the pseudo view synthesis. For the VecTree Compression, we set the codebook size to 8192.
Figure 5: Visual Comparisons for Ablation Study. We visualize the rendered RGB images and the residual map between the ground-truth image, aligned with the experiment ID as shown in Tab. 2. The final model (Exp #7, #8, #9) demonstrates close results to 3D-GS (Exp #1), while the Gaussian Co-adaptation, along with SH distillation, almost completely mitigates the information loss.
4.2 Experimental Results
Quantitative Results
To assess the quality and model size of different methods for novel view synthesis, we summarize the quantitative numbers in Tab. 1, including highly efficient voxel-based NeRFs (Plenoxel[69], Instant-NGP [46]), the compact MLP-based Mip-NeRF360 [5], NeRF with vector quantization [39], and 3D Gaussian Splatting [32]. Specifically, in the Mip-NeRF360 datasets, the NeRF-based methods, benefiting from their compact representation by using MLPs, show competitive accuracy, but the slow inference speed (0.06FPS) makes them impractical for real-world applications. Voxel-based NeRFs partly resolve the rendering efficiency issue, but the FPS still lags behind for practical usage, while Plenoxel requires 2.1GB for storing a large-scale scene. VQ-DVGO optimizes the storage issue, but their pruning and purely vector quantization make it challenging to generalize to complex large-scale scenes. On the other hand, 3D-GS strikes a balance in rendering quality and real-time rendering speed; however, it requires nearly gigabytes for storing a single scene. Our pipeline, LightGaussian, achieves the fastest rendering speed (1.55$\times$) compared to all existing methods, thanks to its pruning of insignificant Gaussians for an efficient rasterization process. Moreover, the compact SH representation and VecTree compression further reduce the model redundancy in 3D Gaussians, shrinking the model size from 727MB to 42.48MB on Mip-NeRF360 datasets, a 17$\times$ reduction ratio. LightGaussian also achieves nearly 2$\times$ faster rendering speed on the Tank & Temple datasets, while reducing the storage from 380MB to 22.43MB.
Qualitative Results
We compare the rendering result of all adopted baselines, with the highlight on intricate details and the background regions in Fig. 4. We can observe Mip-NeRF360, 3D-GS and our LightGaussian achieves the best visual quality, whereas Mip-NeRF360 requires more than 30 seconds to render an image of resolution 1080p, 3D-GS requires 17$\times$ times more disk storage requirement.
4.3 Ablation Studies
We ablate components of our method by separately analyzing the proposed modules. Specifically, we find that Iterative Gaussian Pruning is effective in removing redundant Gaussians, SHs Distillation effectively reduces the SH’s degree, and Vector Quantization efficiently compresses the feature space, leading to a more compact and efficient representation. These individual contributions collectively enhance the overall performance of our framework.
Overall Analysis
We report the performance of the investigation on the proposed modules in Tab. 2 and Fig. 5. We verify that our design of Gaussian Pruning & Recovery is effective in removing redundant Gaussians (Exp #1$\rightarrow$3) with negligible quality degradation while preserving rendering accuracy. This proves that the proposed Global Significance, based on Gaussian attributes, accurately represents the critical aspects. By removing the high-degree SHs and transferring the knowledge to a compact representation (Exp #3 $\rightarrow$5), our method successfully demonstrates the benefits from using soft targets and extra data from view augmentation, results in negligible changes in specular reflection In practical post-processing, the vector quantization on the least important Gaussians (Exp #7), combined with the octree-based lossless compression for its geometric position (Exp #8), showcases the advantage of adopting a hybrid compression.
Figure 6: Visualization of Pruned Gaussians. We show the pruned Gaussians (middle) obtained by applying the proposed Gaussian Prune and Recovery. The residual is visualized by rasterizing the pruned Gaussians.
The Best Significance Criteria?
To discover the most favorable criteria to measure the global significance of each Gaussian, we study the critical elements within each Gaussian, including its hit count with all training rays, Gaussian opacity, and the function-related volume. As shown in Tab. 3, considering all these three attributes by weighting the hit count using opacity and Gaussian volume produces the most accurate rendering quality after zero-shot pruning. A brief Gaussian Co-adaptation (last row) could restore accuracy to its level prior to pruning. We visualize the rasterized images before and after pruning, as well as the pruned Gaussians, in Fig. 6.
SH Distillation
Directly eliminating the high-degree component of SHs leads to an unacceptable performance drop from the full model (Exp #3 $\rightarrow$ 4), where we can also observe a significant loss of specular reflection on surfaces when changing viewpoints. However, by introducing knowledge distillation (Exp #4 $\rightarrow$ 5) from the full model, we can reduce its size and preserve the viewing effects. Additionally, adding pseudo-views during training (Exp #5 $\rightarrow$ 6) further demonstrates the effectiveness of guiding the training of a student model through the teacher model.
VecTree Attribute Compression
We study the effectiveness of applying quantization on Gaussian attributes, as shown in Tab. 4. We find that the Gaussian location is sensitive to a low-bit representation and thus apply lossless Octree-based compression on it. For the rest of the attributes, we first quantize them to half-precision, and applying Vector Quantization (VQ) on all attributes leads to inferior results, but this can be mitigated by using Gaussian Global Significance. The final module, VecTree Compression, performs Octree-based compression on Gaussian location, VQ on the least important SHs, and FP16 on the rest, achieving the best performance balance.
5 Conclusion
We present LightGaussian, a novel framework that converts the heavy point-based representation into a compact format for efficient novel view synthesis. For practical use, LightGaussian explores the use of 3D Gaussians for modeling large-scale scenes and finds an effective way to identify the least important Gaussians grown by densification. In pursuing a compact Spherical Harmonics format, distillation is enhanced by synthesizing pseudo-views to generate more data for knowledge transfer. A VecTree compression post-processing effectively removes further redundancy by using Gaussian significance as an indicator. With these methods, the proposed representation reduces data redundancy by more than 15$\times$, further boosting the FPS to more than 200 FPS, with minimal rendering quality loss.
6 Acknowledgements
We would like to express our gratitude to Yueyu Hu from NYU for the invaluable discussion on our project.
A7 More Technical Details
We detail the procedures of LightGaussian in Algorithm 1. The trained 3D-GS [32] features a Gaussian location with a dimension of 3, Spherical Harmonics coefficients with a dimension of 48, and opacity, rotation, and scale, whose dimensions are 1, 4, and 3, respectively.
Algorithm 1 The overall pipeline of LightGaussian
1:Training view images =$\mathcal{I}=\{\boldsymbol{I}_{i}\in{}^{M}\}_{i=1}^{N}$ and their associated camera poses $\mathcal{P}=\{\boldsymbol{\phi}_{i}\in{}^{3\times 4}\}_{i=1}^{N}$.
2:# Pre-Training 3D-GS [32].
3:# $\boldsymbol{G_{i}}$ with attributes XYZ, SH-3deg, Opacity, Rotation, Scale.
4:#Gaussian Pruning and Recovery.
[…]
9:#Distilling into Compact SHs.
[…]
11:while Few Steps do $\triangleright$ SH Distillation
12:    $\mathcal{\hat{P}}=\text{SampleView}(\mathcal{P})$triangleright Synthesize Pseudo Views
13:    $I_{t}\leftarrow\textsc{\text{Teacher}}(\mathcal{\hat{P}})$triangleright Teacher render
14:    $I_{s}\leftarrow\textsc{\text{Student}}(\mathcal{\hat{P}})$triangleright Student render
15:    $\nabla L\leftarrow\textsc{Loss}(I_{s},I_{t})$
16:    $\mathcal{\hat{G}}\leftarrow\textsc{Adam}(\nabla L)$triangleright Backprop & Step
17:end while
18:#VecTree Quantization.
[…]
21:$\text{XYZ}^{\prime}\leftarrow$ OctreeCompression(XYZ)
22:#Save Model.
23:Save optimized model $\mathcal{\hat{G}}$ to disk.
Specifically, we describe how we calculate the Global Significance Score for each trained Gaussian in Algorithm 2.
Algorithm 2 Global significance calculation for gaussians.
1:# $\mathcal{G}$ contains all gaussians with attributes XYZ, SH-3deg, Opacity, Rotation, Scale.
2:# $\mathcal{P}$ Sample Camera poses from training.
3:function $CalGS$($\mathcal{G},\mathcal{P}$)
4:    $GS\leftarrow 0$triangleright Init Global Significance
5:    for all pixels $\ i\ \ \text{in renderFunc}(\mathcal{P})$ do
6:         $H\leftarrow$ GetHitCount($\mathcal{G},i$) $\triangleright$ Gaussian Hit count
7:         # Eq.5 in the main draft.
8:         $GS\leftarrow GS+H\cdot Opacity^{\intercal}\cdot$ VNorm($Scale$${}^{\intercal}$
9:    end for
10:    return $GS$
11:end function
A8 More Experiment Results
In addition to realistic indoor and outdoor scenes in the Mip-NeRF360 [5] and Tank and Temple datasets [34], we further evaluate our method on the synthetic Blender dataset [44], and provide a scene-wise evaluation on all datasets, accompanied by detailed visualizations.
Results on NeRF-Synthetic 360°(Blender) Dataset.
The synthetic Blender dataset [44] includes eight photo-realistic synthetic objects with ground-truth controlled camera poses and rendered viewpoints (100 for training and 200 for testing). Similar to 3D-GS [32], we start training the model using random initialization. Consequently, we calculate the Global Significance of each Gaussian, work to reduce the SH redundancy, and apply the VecTree Compression (codebook size set at 8192) to the learned representation. Overall comparisons with previous methods are listed in Table A5, where we observe our LightGaussian markedly reduces the average storage size from 52.38MB to 7.89MB, while improving the FPS from 310 to 411 with only a slight rendering quality decrease.
Table A5: Per-scene results on Synthetic-NeRF.
Method Chair Drums Ficus Hotdog Lego Materials Mic Ship Avg.
Size(MB)
3D-GS 94.612 64.163 35.839 39.607 64.910 29.335 34.185 56.400 52.381
Ours 13.785 9.596 5.473 5.994 9.600 4.542 5.252 8.464 7.838
PSNR(dB)
3D-GS 35.436 26.294 35.614 37.848 35.782 30.533 36.585 31.642 33.716
Ours 34.769 26.022 34.484 36.461 34.944 29.341 35.370 30.405 32.725
SSIM
3D-GS 0.987 0.954 0.987 0.985 0.981 0.961 0.992 0.904 0.969
Ours 0.986 0.952 0.985 0.982 0.979 0.954 0.990 0.896 0.965
LPIPS
3D-GS 0.0133 0.0405 0.0121 0.0232 0.0195 0.0404 0.0073 0.111 0.0334
Ours 0.0142 0.0431 0.0137 0.0275 0.0222 0.0461 0.0087 0.121 0.0370
Additional Qualitative Results on Mip-NeRF360.
We provide extra visualizations for 3D-GS [32], LightGaussian (ours), and VQ-DVGO [39], accompanied by the corresponding residual maps from the ground truth. As evidenced in Figure A7 and Figure A8, LightGaussian outperforms VQ-DVGO, which utilizes NeRF as a basic representation. Furthermore, LightGaussian achieves a comparable rendering quality to 3D-GS [32], demonstrating the effectiveness of our proposed compact representation.
Additional Quantitative Results on Mip-NeRF360.
Tables A6, A7, and A8 present the comprehensive error metrics compiled for our evaluation across all real-world scenes (Mip-NeRF360 and Tank and Temple datasets). Our method not only compresses the average model size from 727MB to 42MB, but also consistently demonstrates comparable metrics with 3D-GS on all scenes. LightGaussian additionally shows better rendering quality than Plenoxel, INGP, mip-NeRF360, and VQ-DVGO.
Figure A7: Additional Visual Comparisons on the Mip-NeRF360 Datasets. We present the rendering results from 3D-GS [32], LightGaussian, and VQ-DVGO [39]. The corresponding residual maps highlighting the differences between the rendered images and ground truth (GT) images are also displayed.
Figure A8: Additional Visual Comparisons on the Mip-NeRF360 Datasets. We present the rendering results from 3D-GS [32], LightGaussian, and VQ-DVGO [39]. The corresponding residual maps highlighting the differences between the rendered images and ground truth (GT) images are also displayed.
Table A6: Quantitative Comparison (PSNR) for Mip-NeRF360 and Tank & Temple Scenes.
Methods PSNR
bicycle garden stump room counter kitchen bonsai truck(T&T) train(T&T)
Plenoxels 21.912 23.4947 20.661 27.594 23.624 23.420 24.669 23.221 18.927
INGP-Big 22.171 25.069 23.466 29.690 26.691 29.479 30.685 23.383 20.456
mip-NeRF360 24.305 26.875 26.175 31.467 29.447 31.989 33.397 24.912 19.523
VQ-DVGO 22.089 24.119 23.455 28.423 26.084 25.930 27.985 - -
3D-GS 25.122 27.294 26.783 31.687 29.114 31.618 32.307 24.778 21.449
Ours 24.960 26.735 26.701 31.271 28.113 30.402 31.014 24.561 21.095
Implementation Details of VQ-DVGO.
In the implementation of VQ-DVGO [39], we initially obtain a non-compressed grid model following the default training configuration of DVGO [56]. The pruning quantile $\beta_{p}$ is set to 0.001, the keeping quantile $\beta_{k}$ is set to 0.9999, and the codebook size is configured to 4096. We save the volume density and the non-VQ voxels in the fp16 format without additional quantization. For the joint finetuning process, we have increased the iteration count to 25,000, surpassing the default setting of 10,000 iterations, to maximize the model’s capabilities. All other parameters are aligned with those specified in the original VQ-DVGO paper [39], ensuring a faithful replication of established methodologies.
Table A7: Quantitative Comparison (SSIM) for Mip-NeRF360 and Tank & Temple Scenes.
Methods SSIM
bicycle garden stump room counter kitchen bonsai truck(T&T) train(T&T)
Plenoxels 0.496 0.6063 0.523 0.8417 0.759 0.648 0.814 0.774 0.663
INGP-Big 0.512 0.701 0.594 0.871 0.817 0.858 0.906 0.800 0.689
mip-NeRF360 0.685 0.809 0.745 0.910 0.892 0.917 0.938 0.857 0.660
VQ-DVGO 0.473 0.613 0.564 0.860 0.777 0.763 0.842 - -
3D-GS 0.746 0.856 0.770 0.914 0.914 0.932 0.946 0.863 0.781
Ours 0.738 0.836 0.768 0.926 0.893 0.914 0.933 0.855 0.760
Table A8: Quantitative Comparison (LPIPS) for Mip-NeRF360 and Tank & Temple Scenes.
Methods LPIPS
bicycle garden stump room counter kitchen bonsai truck(T&T) train(T&T)
Plenoxels 0.506 0.3864 0.503 0.4186 0.441 0.447 0.398 0.335 0.422
INGP-Big 0.446 0.257 0.421 0.261 0.306 0.195 0.205 0.249 0.360
mip-NeRF360 0.305 0.171 0.265 0.213 0.207 0.128 0.179 0.159 0.354
VQ-DVGO 0.572 0.404 0.424 0.219 0.244 0.219 0.191 - -
3D-GS 0.245 0.122 0.242 0.200 0.186 0.118 0.184 0.175 0.262
Ours 0.265 0.155 0.261 0.220 0.218 0.147 0.204 0.188 0.296

### v4

@@
 Additionally, these representations are utilized in shadow [46] and cloud rendering applications [41].
-Recently, 3D Gaussian Splatting (3D-GS) [32], combining the idea of point-based rendering and splatting techniques for rendering, achieves real-time speed with comparable rendering quality to the best MLP-based renderer, Mip-NeRF 360 [5].
+Recently, 3D Gaussian Splatting (3D-GS) [32], combining the idea of point-based rendering and splatting techniques [72] for rendering, achieves real-time speed with comparable rendering quality to the best MLP-based renderer, Mip-NeRF 360 [5].
 Although promising, the heavy storage requirement to store all attributes attached to the Gaussians often requires gigabyte-level disk space for saving a single unbounded scene, and millions of Gaussians hinder the rendering efficiency of 3D-GS.
@@
 Here, $c_{i}$, $\alpha_{i}$ represents the color and opacity of this point computed by a Gaussian with covariance $\boldsymbol{\Sigma}$ multiplied by an optimizable per-point opacity and SH color coefficients.
-In summary, each Gaussian point is characterized by attributes including: position $\boldsymbol{X}\in\mathbb{R}^{3}$, color defined by spherical harmonics coefficients $\boldsymbol{C}\in\mathbb{R}^{(k+1)^{2}}\times 3$ (where $k$ represents the degrees of freedom), opacity $\alpha\in\mathbb{R}$, rotation factor $\boldsymbol{r}\in\mathbb{R}^{4}$, and scaling factor $\boldsymbol{s}\in\mathbb{R}^{3}$.
+In summary, each Gaussian point is characterized by attributes including: position $\boldsymbol{X}\in\mathbb{R}^{3}$, color defined by spherical harmonics coefficients $\boldsymbol{C}\in\mathbb{R}^{(k+1)^{2}}\times 3$ (where $k$ represents the degrees of freedom), opacity $\alpha\in\mathbb{R}$, rotation factor $\mathbf{R}\in\mathbb{R}^{4}$, and scaling factor $\mathbf{S}\in\mathbb{R}^{3}$.
 3.2 Gaussian Pruning & Recovery
@@
 Such an explicit point-based representation with a large number of Gaussians requires an extremely significant storage overhead.
+Drawing inspiration from the success of representative neural network pruning techniques [26], which eliminate less impactful neurons without compromising the network’s overall performance, we tailor a general pruning paradigm for point-based representation to reduce the over-parameterized point number in a manner that preserves the original accuracy.
+Therefore, identifying the most representative redundant Gaussians with recoverable accuracy is a crucial step in our approach.
 However, pruning the Gaussians based on simplistic criteria (e.g., point opacity) can lead to a substantial degradation in modeling performance, especially where the intricate scene structure may be eliminated, as is demonstrated in Fig.
@@
 This has inspired us to find better criteria to measure global significance in terms of rendering quality.
-Drawing inspiration from the success of representative neural network pruning techniques [26], which eliminate less impactful neurons without compromising the network’s overall performance, we tailor a general pruning paradigm for point-based representation to reduce the over-parameterized point number in a manner that preserves the original accuracy.
-Therefore, identifying the most representative redundant Gaussians with recoverable accuracy is a crucial step in our approach.
+The accumulated PDF is equal to 1.
 Global Significance Calculation
 Simply relying on Gaussian opacity as a significance criterion leads to suboptimal Gaussian pruning, prompting the need for a more effective formula.
-Inspired by Equation 3, these 3D Gaussians can be rasterized by projecting them onto a specific camera viewpoint for image rendering.
-The significance of each Gaussian can then be quantified based on its contribution to each pixel across all training views, with similar principle in magnitude network pruning [38].
-Consequently, we iterate over all training pixels to calculate the hit count of each Gaussian, factoring in both Gaussian opacity and volume:
-$\text{GS}_{\text{j}}$ $=\sum_{i=1}^{MHW}\mathbbm{1}(\text{G}(\boldsymbol{X}_{j}),r_{i})\cdot\sigma_{j}\cdot\gamma_{j}(\boldsymbol{\Sigma}),$ (4)
+Inspired by Eq 3, these 3D Gaussians can be rasterized by projecting them onto a specific camera viewpoint for image rendering.
+The initial significance score of each Gaussian can then be quantified based on its contribution to each pixel (a.k.a rays, $r_{i}$) across all training views using the criteria $\mathbbm{1}(\text{G}(\boldsymbol{X}_{j}),r_{i})$.
+Consequently, we iterate over all training pixels to calculate the hit count of each Gaussian.
+The score is further refined by the adjusted 3D Gaussian’s volume $\large{\gamma}$$(\boldsymbol{\Sigma}_{j})$ and opacity $\sigma_{j}$.
+The volume calculation equation is V${(\boldsymbol{\Sigma}_{j})}=\frac{4}{3}\pi abc$, where $abc$ are the 3 dimensions of Scale ($\mathbf{S}$).
+We implement this in PyTorch as torch.prod(gaussians.scaling, dim=1) without considering the constant value $\frac{4}{3}\pi$.
+Overall, the global significant score can be summarized as:
+$\text{GS}_{\text{j}}$ $=\sum_{i=1}^{MHW}\mathbbm{1}(\text{G}(\boldsymbol{X}_{j}),r_{i})\cdot\sigma_{j}\cdot{\large{\gamma}}(\boldsymbol{\Sigma}_{j}),$ (4)
 where $j$ indicates the Gaussian index, and $M$, $H$, and $W$ represent the number of training views, image height, and width, respectively.
 $\mathbbm{1}$ is the indicator function that determines whether a Gaussian intersects with a given ray.
-The use of absolute Gaussian volume tends to exaggerate the importance of background Gaussians, leading to the immoderate pruning of Gaussians that model intricate geometry.
+However, the use of Gaussian volume tends to exaggerate the importance of background Gaussians, leading to the immoderate pruning of Gaussians that model intricate geometry.
 Therefore, we propose a more adaptive way to measure the dimension of its volume:
@@
 Stripping away too many Gaussians can lead to a noticeable degradation in the model’s accuracy and visual fidelity since incomplete scene coverage.
-To mitigate this, a phase of short Gaussian co-adaptation is essential, which involves a joint adjustment of the Gaussians’ attributes, allowing them to adapt and compensate for the loss incurred by pruning.
+To mitigate this, we propose to utilize photometric loss on original training views to optimize (adapt) the remaining Gaussians, which jointly adjusts their attributes over 5k steps, without further densification, which can compensate for the loss incurred by pruning.
 3.3 Distilling into Compact SHs
@@
 Reducing the degree of SH, while beneficial for decreasing disk space usage, leads to a noticeable loss of surface ‘shininess’, particularly affecting specular reflection variance when the viewpoint is altered.
-To achieve a balance between disk storage efficiency and scene reflectance quality, we propose transferring knowledge from well-trained high-degree SHs to their compact counterparts (lower-degree) via data distillation:
-$\mathcal{L}_{\text{distill}}(\text{G}_{\text{SH2-deg}})=\frac{1}{N}\sum_{i=1}^{N}\left\lVert\boldsymbol{C}_{\text{teacher}}-\boldsymbol{C}_{\text{student}}\right\rVert_{2}^{2}.$ (6)
+To achieve a balance between model size and scene reflectance quality, we propose transferring knowledge from well-trained high-degree SHs to their compact counterparts (lower-degree) via distillation.
+Specifically, we distill the teacher model (with full-degree SHs, and is frozen) to the student model (with truncated lower-degree SHs).
+The supervisions are from the difference of the predicted pixel intensities $\boldsymbol{C}$ from the two models :
+$\mathcal{L}_{\text{distill}}=\frac{1}{HW}\sum_{i=1}^{HW}\left\lVert\boldsymbol{C}_{\text{teacher}}(r_{i})-\boldsymbol{C}_{\text{student}}(r_{i})\right\rVert_{2}^{2}.$ (6)
 Simply reintroducing these view-dependent visual effects to the network does not significantly enhance its knowledge.
@@
 Although a high compression rate can be reached, Gaussian attributes cannot be easily quantized in the same fashion.
-The reason is that each attribute of the explicit representation has its own physical meaning, especially when representing large-scale 3D scenes, and attributes such as opacity, rotation, and scale are more sensitive to such discretized representation, with a dramatic drop in accuracy.
+The reason is that each attribute of the explicit representation has its own physical meaning and complex correlation (e.g., the rotation and scaling formulate the covariance matrix), and we empirically find that certain attributes such as opacity, rotation, and scale are more sensitive to such discretized representation, with a dramatic drop in accuracy.
 Therefore, we propose two ways to mitigate this gap: first, we reutilize the significance score from Sec.
@@
 Vector Quantization
-The codebook under vector quantization (VQ) is initialized with $K$ codes within it, denoted as $\boldsymbol{C}={\boldsymbol{c}_{1},\boldsymbol{c}_{2},\boldsymbol{c}_{3},\boldsymbol{c}_{K}}$ to represent the Gaussian sets $\boldsymbol{G}={\boldsymbol{g}_{1},\boldsymbol{g}_{2},\boldsymbol{g}_{3},\boldsymbol{g}_{N}}$ where $K\ll N$.
-To assign each code to the Gaussian feature, we randomly select a batch of Gaussians at each iteration and calculate the Euclidean distance between the selected Gaussian and each code vector in the codebook to determine which code the Gaussian is associated with.
-Then, a significance-weighted codebook optimization is applied: $\mathbf{c}_{i}:=\lambda\cdot\mathbf{c}_{i}+(1-\lambda)\cdot\sum{\mathbf{g}_{j}\in\mathcal{R}(\mathbf{c}_{k})}\text{GS}_{j}\cdot\mathbf{g}_{j}$, where $\mathbf{g}_{j}\in\mathcal{R}(\mathbf{c}_{i})$ denotes the $j$-th Gaussian assigned to the $i$-th code, and $\lambda$ controls the contribution from the moving average.
-We only perform excessive VQ on the least 60% of Gaussians, since we are representing complex, large-scale scenes.
+VQ partitions the Gaussians $\mathcal{G}=\{\mathbf{g}_{1},\mathbf{g}_{2},\ldots,\mathbf{g}_{N}\}$ to the K codes in the codebook $\mathcal{C}=\{\mathbf{c}_{1},\mathbf{c}_{2},\ldots,\mathbf{c}_{K}\}$, where each $\mathbf{g}_{j},\mathbf{c}_{k}\in\mathbb{R}^{d}$ and K $\ll$ N.
+$d$ means attributes’ dimension.
+Specifically, $\mathcal{C}$ is initialized via K-means.
+Then, VQ iteratively samples a batch of $\mathcal{G}$, associates them to the closest codes by euclidean distance, and updates each $\mathbf{c}_{k}$ via moving average rule: $\mathbf{c}_{k}=0.8\cdot\mathbf{c}_{k}+(1-0.8)\cdot 1/\mathbf{T}_{k}\cdot\sum_{\mathbf{g}_{j}\in\mathcal{R}(\mathbf{c}_{k})}\mathbf{GS}_{j}\cdot\mathbf{g}_{j}$, where $\mathbf{T}_{k}=\sum_{\mathbf{g}_{j}\in\mathcal{R}(\mathbf{c}_{k})}{\mathbf{GS}_{j}}$ is the total significance score assigned to the code vector $\mathbf{c}_{k}$, $\mathcal{R}(\mathbf{c}_{k})$ is the set of Gaussians associated to the $k$-th code.
 Methods Mip-NeRF 360 Datasets Tank and Temple Datasets
@@
 Octree-based Compression
-Unstructured and high-precision 3D points pose a challenge for efficient storage, especially for large-scale scenes.
-Applying lossy compression to Gaussian locations shows significant rendering degradation, as the Gaussian location is sensitive to the subsequent rasterization accuracy.
-Therefore, drawing from the Point Cloud Compression literature [53], which applies lossless encoding on the Gaussian location, utilizing the octree geometry codec in G-PCC [45] consumes only a very small amount of bits.
-Specifically, let $(X_{i}=(x_{i},y_{i},z_{i}))_{i=1\ldots N}$ be the set of 3D positions associated with the points of the input point cloud.
-The L-PCC encoder [53] computes the quantized positions $(\hat{X}_{i})_{i=1...N}$ as follows: $\hat{X}_{i}=\left\lfloor(X_{i}-X_{shift})\times s\right\rfloor,$ where $X_{shift}$ and $s$ are pre-defined parameters and we set all octree-related parameters the same as PCGCv2 [62].
-For the decoding stage, the reconstructed positions $(\hat{X}_{i})_{i=1}^{N}$ are generated by applying the following inverse quantization process: $\hat{X}_{i}=\frac{\hat{X}i}{s}+X_{shift}$.
+To deal with the unstructured and high-precision 3D points location, we avoid the lossy compression which is sensitive to the subsequent rasterization accuracy.
+Drawing inspiration from the Point Cloud Compression literature [52], which applies lossless encoding on the Gaussian location, utilizing the octree geometry codec in G-PCC [44] consumes only a very small amount of bits.
 4 Experiments
@@
 Qualitative Results
-We compare the rendering result of all adopted baselines, with the highlight on intricate details and the background regions in Fig.
-4.
-We can observe Mip-NeRF360, 3D-GS and our LightGaussian achieves the best visual quality, whereas Mip-NeRF360 requires more than 30 seconds to render an image of resolution 1080p, 3D-GS requires 17$\times$ times more disk storage requirement.
+We compared the rendering results between 3D-GS and ours, with an emphasis on intricate details and background regions as illustrated in Fig.
+4, to highlight the differences.
+It is observable that 3D-GS and our LightGaussian achieve similar visual quality, even in thin structures within the scenes (the first and third rows).
+The Gaussian compression method slightly affects the specular effect on the surface (see the second and the fourth rows).
 4.3 Ablation Studies
@@
 The final module, VecTree Compression, performs Octree-based compression on Gaussian location, VQ on the least important SHs, and FP16 on the rest, achieving the best performance balance.
+Degradation & Speed vs. Compression Ratio?
+To strike a balance among model compression ratio and the reduced model size as well as the rendering speed, we study the effect of various ratios for Gaussian Pruning & Recovery, SHs Distilling, and VecTree Compression.
+As can be seen in the Fig 7 and Tab 5.
+The ratios selected represent a balanced point to optimize the trade-off between model size, FPS, and rendering quality (see Tab 5).
+Figure 7: Visualization of The effect from different prune and VQ ratio.
+Exp ID Model setting Model Size $\downarrow$ FPS $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
+[1] 3D-GS 353.27 MB 192 31.687 0.927 0.200
+[2] 3D-GS (w/ 40% pruned Gaussian) 211.96 MB 245 31.930 0.928 0.198
+[3] [2] + 2-degree SH compactness 140.16 MB 238 30.625 0.917 0.208
+[4] [3] + proposed distillation 140.16 MB 243 31.754 0.926 0.201
+[5] [2] + 1-degree SH compactness 88.89 MB 249 29.173 0.900 0.223
+[6] [5] + proposed distillation 88.89 MB 244 31.440 0.923 0.205
+[7] [4] + 55% vector quantization 42.92MB 244 31.626 0.924 0.204
+Table 5: Ablation under a mild compression ratio: With a Gaussian pruning ratio of 40% and VQ ratio of 55%, LightGaussian nearly preserves the rendering quality (SSIM decrease from 0.927 to 0.924), while achieving an $\sim$87reduction in model size and enhancing the FPS from 192 to 244.
 5 Conclusion
@@
 Additional Quantitative Results on Mip-NeRF360.
-Tables A6, A7, and A8 present the comprehensive error metrics compiled for our evaluation across all real-world scenes (Mip-NeRF360 and Tank and Temple datasets).
+Tables A7, A8, and A9 present the comprehensive error metrics compiled for our evaluation across all real-world scenes (Mip-NeRF360 and Tank and Temple datasets).
 Our method not only compresses the average model size from 727MB to 42MB, but also consistently demonstrates comparable metrics with 3D-GS on all scenes.

### v5

@@
 Drawing inspiration from the concept of Network Pruning, LightGaussian identifies Gaussians that are insignificant in contributing to the scene reconstruction and adopts a pruning and recovery process, effectively reducing redundancy in Gaussian counts while preserving visual effects.
-Additionally, LightGaussian employs distillation and pseudo-view augmentation to distill spherical harmonics to a lower degree, allowing knowledge transfer to more compact representations while maintaining scene appearance.
-Furthermore, we propose a hybrid scheme, VecTree Quantization, to quantize all attributes, resulting in lower bitwidth representations with minimal accuracy losses.
-In summary, LightGaussian achieves an averaged compression rate over 15$\times$ while boosting the FPS from 139 to 215, enabling an efficient representation of complex scenes on Mip-NeRF 360, Tank & Temple datasets.
-Project website: https://lightgaussian.github.io/
+Additionally, LightGaussian employs knowledge distillation and pseudo-view augmentation to transfer spherical harmonics efficients to a lower degree, allowing knowledge convert to more compact representations while maintaining scene appearance.
+Furthermore, we propose a Gaussian Vector Quantization based on Gaussian global significance, to quantize all redundant attributes, resulting in lower bitwidth representations with minimal accuracy losses.
+In summary, LightGaussian achieves an averaged compression rate over 15$\times$ while boosting the FPS from 119 to 209, enabling an efficient representation of complex scenes on Mip-NeRF 360, Tank & Temple datasets.
 Figure 1: Compressibility and Rendering Speed.
@@
 LightGaussian effectively prunes redundant Gaussians while preserving visual fidelity (on the left).
-Consequently, it reduces the average storage from 727MB to 42MB and improves the FPS from 139 to 215.
-11footnotetext: Equal Contribution.
+It reduces the average storage from 724MB to 42MB and improves the FPS from 119 to 209.
 1 Introduction
@@
 Subsequent efforts along this trend introduce voxel-based implicit representation [57, 62, 54, 67, 10], hash grid [46], computational parallelism [51], or neural light field representation [61] to speed up the rendering.
-However, they either require task-specific design or face a trade-off between rendering quality and efficiency, making them difficult to generalize to large-scale scenarios with practical speed.
+However, they either require specific design or face a trade-off between rendering quality and efficiency, making them difficult to generalize to large-scale scenarios with practical speed.
 Recent progress in point-based 3D Gaussian Splatting (3D-GS) [32] has brought photo-realistic rendering quality to the real-time level, even for complex scenes.
@@
 Attributes are attached to Gaussians to model the scene’s geometry and appearance.
-However, the significant number of Gaussians, as well as the high-degree Spherical Harmonics (SH) coefficients used for modeling scene reflection, contribute to an over-parametrized representation when fitting to the scenes.
+However, the significant number of Gaussians after its optimization, as well as the high-degree Spherical Harmonics (SH) coefficients used for modeling scene reflection, contribute to an over-parametrized representation when fitting to the scenes.
 To minimize the required Gaussian number, we propose proper criteria to measure the global significance of each 3D Gaussian, in the context of its contribution for view synthesis.
-Gaussians with low impact on visual quality will be identified and pruned, followed by short recovery steps to be applied.
-Spherical harmonics (SH) coefficients, as the majority of the data, are used for modeling the view-dependent color.
-Compressing it directly by shrinking the higher degree harms the reflectance effect, and we propose a general distillation step enhanced by pseudo-view augmentation to harmlessly transfer the knowledge into a compact level.
-VecTree Quantization step adaptively picks “just some amount” of distinct point attributes based on the global significance, further reducing the required bitwidth of the original format.
+Gaussians that have minimal impact on visual quality are identified and pruned.
+This is followed by streamlined recovery steps that allow for rapid optimization, named as Gaussian Pruning and Recovery.
+Spherical harmonics (SH) coefficients, constituting a significant portion of the data, are employed to model view-dependent color and to capture the interaction of light with surfaces from diverse angles.
+However, utilizing higher-degree SH coefficients presents a trade-off between accuracy and computational efficiency.
+While higher-degree coefficients offer more detailed lighting information, enhancing scene realism, they also escalate computational complexity and storage demands.
+To mitigate this, we propose a SH Distillation step complemented by pseudo-view augmentation.
+This approach is designed to efficiently condense the information from higher-degree coefficients into a more compact form, ensuring that quality is preserved without inflating computational requirements.
+Besides focusing on reducing redundancy in spatial (Gaussian count) and lighting (SHs) representations, we additionally introduce a Vector Quantization step.
+This process adaptively selects a specific quantity of distinct point attributes, determined by their global significance, which further reduces the required bitwidth from the original format.
 In summary, our proposed framework, LightGaussian, efficiently reduces the Gaussian count (e.g., from 1.49M to 575K, Fig.
-1, left), significantly reducing storage requirements from 727MB to 42MB, with minimal rendering quality decrease ($\downarrow$ 0.013 in SSIM) on the Mip-NeRF 360 datasets (Fig.
+1, left), significantly reducing storage requirements from 724MB to 42MB, with minimal rendering quality decrease ($\downarrow$ 0.005 in SSIM) on the Mip-NeRF 360 dataset (Fig.
 1, right).
@@
 2 Related Works
-Figure 2: The overall pipeline of LightGaussian.
-3D Gaussians are optimized from multi-view images and SfM points.
-LightGaussian first calculates the global significance for each Gaussian based on training observations; Gaussians with the least significance are pruned.
-A subsequent distillation with synthesized pseudo-views is introduced to transfer the SH into a compact format.
-VecTree quantization, consisting of lossy codebook quantization and lossless octree quantization, is further introduced for compressing the representation.
-Efficient 3D Scene Representations for NVS
-Neural radiance field (NeRF) [43] uses a multi-layer perceptron (MLP) to represent a scene, and this compact representation has brought view synthesis quality to a new stage.
+Efficient 3D Scene Representations for NVS.
+Neural radiance field
+(NeRF) [45] propose to use multi-layer perceptron (MLPs) to represent a scene, and this compact representation has brought view synthesis quality to a new stage.
 However, NeRF suffers from extremely slow inference challenges.
@@
 Additionally, these representations are utilized in shadow [49] and cloud rendering applications [43].
-Recently, 3D Gaussian Splatting (3D-GS) [32], combining the idea of point-based rendering and splatting techniques [72] for rendering, achieves real-time speed with comparable rendering quality to the best MLP-based renderer, Mip-NeRF 360 [5].
+Recently, Pulsar [37] has shown its efficient sphere rasterization, 3D Gaussian Splatting (3D-GS) [32] utilizes anisotropic Gaussians [72] and introduces a tile-based sorting for rendering, achieves real-time speed with comparable rendering quality to the MLP-based ones(e.g., Mip-NeRF 360 [5]).
 Although promising, the heavy storage requirement to store all attributes attached to the Gaussians often requires gigabyte-level disk space for saving a single unbounded scene, and millions of Gaussians hinder the rendering efficiency of 3D-GS.
-Model Pruning and Vector Quantization
+Concurrent works adopt region-based vector quantization [48], K-means based codebooks [47], or learned binary masks for each Gaussian, replacing the SHs with grid-based neural networks [39] within a modified training framework to reduce model size.
+Model Pruning and Vector Quantization.
 Model pruning involves reducing the complexity of a neural network by eliminating non-significant parameters to maintain a balance between performance and resource utilization.
@@
 In this vein, vector quantization has been widely adopted in image synthesis [16], text-to-image generation [24], and novel view synthesis [68, 40, 25].
-Knowledge Distillation
+Knowledge Distillation.
 Knowledge Distillation (KD) is widely adopted in various machine learning tasks [55, 53, 63, 61].
@@
 Our work falls into the category of utilizing knowledge distillation for model compression and leverages novel view renderings as the bridge to connect the teacher and student models.
-3 Method
-Overview
+3 Methods
+Figure 2: The overall pipeline of LightGaussian.
+3D Gaussians are optimized from multi-view images and SfM points.
+LightGaussian first calculates the global significance for each Gaussian on training data, where Gaussians with the least significance are pruned.
+A subsequent distillation with synthesized pseudo-views is introduced to transfer the SH into a compact format.
+Vector quantization, consisting of codebook initialization and assignment, is further introduced for reducing the model bandwidth.
+Overview.
 The overview of LightGaussian is illustrated in Fig.
 2.
-The 3D-GS model is trained using multi-view images and is initially initialized from SfM point clouds.
+The 3D-GS model is trained using multi-view images and is initialized from SfM point clouds.
 By expanding the sparse points to millions of Gaussians, the scene is well-represented.
 Then, the 3D-GS undergoes processing within our pipeline to transform it into a more compact format.
-This involves utilizing Gaussian Prune and Recovery to reduce the number of Gaussians, SH Distillation to remove redundant SHs while preserving the modeled specular light, and VecTree Quantization to store Gaussians at a lower bit-width.
+This involves utilizing Gaussian Prune and Recovery to reduce the number of Gaussians, SH Distillation to remove redundant SHs while preserving the essential lighting information, and Vector Quantization to store Gaussians at a lower bit-width.
 3.1 Background: 3D Gaussian Splatting
 3D Gaussian Splatting (3D-GS) [32] is an explicit point-based 3D scene representation, utilizing Gaussians with various attributes to model the scene.
-When representing a complex real-world scene, 3D-GS is initialized from an SfM sparse point cloud, and Gaussian Densifications are applied to increase the Gaussian counts that are used for handling small-scale geometry insufficiently covered.
+When representing a complex real-world scene, 3D-GS is initialized from a sparse point cloud generated by SfM, and Gaussian Densification is applied to increase the Gaussian counts that are used for handling small-scale geometry insufficiently covered and over reconstruction.
 Formally, each Gaussian is characterized by a covariance matrix $\boldsymbol{\Sigma}$ and a center point $\boldsymbol{X}$, which is referred to as the mean value of the Gaussian:
@@
 When rendering 2D images from the 3D Gaussians, the technique of splatting [66, 36] is employed for the Gaussians within the camera planes.
-With a viewing transform denoted as $\boldsymbol{W}$ and the Jacobian of the affine approximation of the projective transformation represented by $\boldsymbol{J}$, the covariance matrix $\boldsymbol{\Sigma}^{\prime}$ in camera coordinates can be computed as follows:
-$\boldsymbol{\Sigma}^{\prime}=\boldsymbol{JW}\boldsymbol{\Sigma}\boldsymbol{W}^{T}\boldsymbol{J}^{T}.$ (2)
+With a viewing transform denoted as $\boldsymbol{W}$ and the Jacobian of the affine approximation of the projective transformation represented by $\boldsymbol{J}$, the covariance matrix $\boldsymbol{\Sigma}^{\prime}$ in camera coordinates can be computed as $\boldsymbol{\Sigma}^{\prime}=\boldsymbol{JW}\boldsymbol{\Sigma}\boldsymbol{W}^{T}\boldsymbol{J}^{T}$.
 Specifically, for each pixel, the color and opacity of all the Gaussians are computed using the Gaussian’s representation Eq.
@@
 The blending of $N$ ordered points that overlap the pixel is given by the formula:
-$\boldsymbol{C}=\sum_{i\in N}\boldsymbol{c_{i}}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{i}).$ (3)
-Here, $c_{i}$, $\alpha_{i}$ represents the color and opacity of this point computed by a Gaussian with covariance $\boldsymbol{\Sigma}$ multiplied by an optimizable per-point opacity and SH color coefficients.
-In summary, each Gaussian point is characterized by attributes including: position $\boldsymbol{X}\in\mathbb{R}^{3}$, color defined by spherical harmonics coefficients $\boldsymbol{C}\in\mathbb{R}^{(k+1)^{2}}\times 3$ (where $k$ represents the degrees of freedom), opacity $\alpha\in\mathbb{R}$, rotation factor $\mathbf{R}\in\mathbb{R}^{4}$, and scaling factor $\mathbf{S}\in\mathbb{R}^{3}$.
+$\boldsymbol{C}=\sum_{i\in N}{c_{i}}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{i}).$ (2)
+Here, $c_{i}$, $\alpha_{i}$ represents the view-dependent color and opacity, calculated from a 2D Gaussian with covariance $\boldsymbol{\Sigma}$ multiplied by an optimizable per-3D Gaussian’s opacity.
+In summary, each Gaussian point is characterized by attributes including: position $\boldsymbol{X}\in\mathbb{R}^{3}$, color defined by spherical harmonics coefficients $\text{SH}\in\mathbb{R}^{(k+1)^{2}}\times 3$ (where $k$ represents the degrees of freedom), opacity $\alpha\in\mathbb{R}$, rotation factor $\mathbf{R}\in\mathbb{R}^{4}$, and scaling factor $\mathbf{S}\in\mathbb{R}^{3}$.
 3.2 Gaussian Pruning & Recovery
@@
 Such an explicit point-based representation with a large number of Gaussians requires an extremely significant storage overhead.
-Drawing inspiration from the success of representative neural network pruning techniques [26], which eliminate less impactful neurons without compromising the network’s overall performance, we tailor a general pruning paradigm for point-based representation to reduce the over-parameterized point number in a manner that preserves the original accuracy.
+Drawing inspiration from the success of representative neural network pruning techniques [26], which eliminates less impactful neurons without compromising the network’s overall performance, we tailor a general pruning paradigm for 3D Gaussian Splatting representation to reduce the over-parameterized point number in a manner that preserves the original accuracy.
 Therefore, identifying the most representative redundant Gaussians with recoverable accuracy is a crucial step in our approach.
@@
 This has inspired us to find better criteria to measure global significance in terms of rendering quality.
-The accumulated PDF is equal to 1.
-Global Significance Calculation
-Simply relying on Gaussian opacity as a significance criterion leads to suboptimal Gaussian pruning, prompting the need for a more effective formula.
-Inspired by Eq 3, these 3D Gaussians can be rasterized by projecting them onto a specific camera viewpoint for image rendering.
-The initial significance score of each Gaussian can then be quantified based on its contribution to each pixel (a.k.a rays, $r_{i}$) across all training views using the criteria $\mathbbm{1}(\text{G}(\boldsymbol{X}_{j}),r_{i})$.
+The accumulated Probability Density Function(PDF) is equal to 1.
+Global Significance Calculation.
+Rather than merely depending on Gaussian opacity for assessing significance, we are inspired by Eq.
+2.
+In this context, 3D Gaussians within the view frustum are evaluated and rasterized by projecting them onto a camera viewpoint for rendering.
+The initial significance score of each Gaussian (denoted as $\text{G}(\boldsymbol{X}_{j})$) can then be quantified based on its contribution to each pixel (ray, $r_{i}$) across all training views using the criteria $\mathbbm{1}(\text{G}(\boldsymbol{X}_{j}),r_{i})$ whether they intersect or not.
 Consequently, we iterate over all training pixels to calculate the hit count of each Gaussian.
-The score is further refined by the adjusted 3D Gaussian’s volume $\large{\gamma}$$(\boldsymbol{\Sigma}_{j})$ and opacity $\sigma_{j}$.
-The volume calculation equation is V${(\boldsymbol{\Sigma}_{j})}=\frac{4}{3}\pi abc$, where $abc$ are the 3 dimensions of Scale ($\mathbf{S}$).
-We implement this in PyTorch as torch.prod(gaussians.scaling, dim=1) without considering the constant value $\frac{4}{3}\pi$.
-Overall, the global significant score can be summarized as:
+The score is further refined by the adjusted 3D Gaussian’s volume $\large{\gamma}$$(\boldsymbol{\Sigma}_{j})$ and 3D Gaussian’s opacity $\sigma_{j}$, as they all contribute for the rendering formula.
+The volume calculation equation of the $j$-th Guassian is V${(\boldsymbol{\Sigma}_{j})}=\frac{4}{3}\pi abc$, where $abc$ are the 3 dimensions of Scale ($\mathbf{S}$).
+Finally, the global significant score can be summarized as:
 $\text{GS}_{\text{j}}$ $=\sum_{i=1}^{MHW}\mathbbm{1}(\text{G}(\boldsymbol{X}_{j}),r_{i})\cdot\sigma_{j}\cdot{\large{\gamma}}(\boldsymbol{\Sigma}_{j}),$ (3)
-where $j$ indicates the Gaussian index, and $M$, $H$, and $W$ represent the number of training views, image height, and width, respectively.
+where $j$ is the Gaussian index, $i$ means a pixel, $M$, $H$, and $W$ represents the number of training views, image height, and width, respectively.
 $\mathbbm{1}$ is the indicator function that determines whether a Gaussian intersects with a given ray.
@@
 $\text{V}_{\text{norm}}$ $=\text{min}\left(\text{max}\left(\frac{\text{V}(\boldsymbol{\Sigma})}{\text{V}_{\text{max90}}},0\right),1\right).$
-Here, the volume is firstly normalized by the 90% largest of all sorted Gaussians, clipping the range between 0 and 1, and $\beta$ is introduced to provide additional flexibility.
-Gaussian Co-adaptation
-We rank all the Gaussians based on the computed global significance scores, and provide a quantitative basis for pruning Gaussians with the lowest scores.
-It is worth noting that an overly aggressive pruning ratio can be detrimental to performance.
-Stripping away too many Gaussians can lead to a noticeable degradation in the model’s accuracy and visual fidelity since incomplete scene coverage.
-To mitigate this, we propose to utilize photometric loss on original training views to optimize (adapt) the remaining Gaussians, which jointly adjusts their attributes over 5k steps, without further densification, which can compensate for the loss incurred by pruning.
+Here, the calculated Gaussian volume is firstly normalized by the 90% largest of all sorted Gaussians, clipping the range between 0 and 1, to avoid excessive floating Gaussians derived from vanilla 3D-GS.
+The $\beta$ is introduced to provide additional flexibility.
+Gaussian Co-adaptation.
+We rank all the Gaussians based on the computed global significance scores, which provides a quantitative basis for pruning Gaussians with lower scores.
+The remaining ones are optimized (adapted) to jointly adjust their attributes, without further densification, to compensate for the slight loss incurred by pruning.
+Note that, we utilize photometric loss on original training views with 5,000 iterations which is aligned with 3D-GS training.
 3.3 Distilling into Compact SHs
-In the uncompressed Gaussian Splat data, a significant portion is comprised of Spherical Harmonics (SH) coefficients, requiring (45+3) floating-point values per splat, which represent 81.3 percent of the total attribute volume.
-Reducing the degree of SH, while beneficial for decreasing disk space usage, leads to a noticeable loss of surface ‘shininess’, particularly affecting specular reflection variance when the viewpoint is altered.
-To achieve a balance between model size and scene reflectance quality, we propose transferring knowledge from well-trained high-degree SHs to their compact counterparts (lower-degree) via distillation.
-Specifically, we distill the teacher model (with full-degree SHs, and is frozen) to the student model (with truncated lower-degree SHs).
-The supervisions are from the difference of the predicted pixel intensities $\boldsymbol{C}$ from the two models :
+In the uncompressed Gaussian Splat data, a considerable portion is occupied by Spherical Harmonics (SH) coefficients, necessitating (45+3) floating-point values per splat, which account for 81.3 percent of the total attribute volume.
+While directly diminishing the degree of SH is advantageous for reducing disk space consumption, this action results in a significant reduction in surface “shininess”, notably impacting the variance in specular reflections as the viewpoint changes.
+To strike a balance between model size and scene modeling quality, we suggest transferring knowledge from well-trained high-degree SHs to their more compact, lower-degree counterparts through a distillation process.
+Specifically, we distill the knowledge from the teacher model, which employs full-degree SHs and remains static, to the student model, characterized by truncated lower-degree SHs.
+The supervision is based on the differences in the predicted pixel intensities, denoted as $\boldsymbol{C}$, between the two models:
 $\mathcal{L}_{\text{distill}}=\frac{1}{HW}\sum_{i=1}^{HW}\left\lVert\boldsymbol{C}_{\text{teacher}}(r_{i})-\boldsymbol{C}_{\text{student}}(r_{i})\right\rVert_{2}^{2}.$ (5)
@@
 This involves training the Gaussians not only to represent known views but also to learn behaviors from unseen (pseudo) views modeled by the teacher model, thus broadening their representational capacity.
-Synthesize Pseudo Views
-The specular reflection rays are reflected off the surface of the light source when we move the viewpoints.
-We can augment the training views by sampling additional viewpoints that can reflect such reflectance.
-As we know the camera locations and viewing directions of all training views, we randomly sample pseudo views around each of the training views, obeying a Gaussian distribution, while fixing the camera view directions:
+Synthesize Pseudo Views.
+Specular reflection rays are reflected off the surface of an object when the viewpoints are altered.
+To enrich the training data, we can introduce additional viewpoints through sampling, which are capable of capturing such reflectance characteristics.
+Given our knowledge of the camera locations and viewing directions for all training views, we strategically sample pseudo views around each training view, adhering to a Gaussian distribution, while maintaining consistent camera viewing directions:
 $\mathbf{t}_{\text{pseudo}}=\mathbf{t}_{\text{train}}+\mathcal{N}(0,\sigma^{2}),$ (6)
@@
 $\mathcal{N}$ denotes a Gaussian distribution with mean 0 and variance $\sigma^{2}$, which is added to the original position to generate the new position.
-3.4 VecTree Attribute Compression
-Vector quantization is well-explored in voxel-based NeRF representations [38, 66, 56, 70].
-Although a high compression rate can be reached, Gaussian attributes cannot be easily quantized in the same fashion.
-The reason is that each attribute of the explicit representation has its own physical meaning and complex correlation (e.g., the rotation and scaling formulate the covariance matrix), and we empirically find that certain attributes such as opacity, rotation, and scale are more sensitive to such discretized representation, with a dramatic drop in accuracy.
-Therefore, we propose two ways to mitigate this gap: first, we reutilize the significance score from Sec.
-3.2 and perform excessive vector quantization on the SHs of the least important Gaussians.
-For the Gaussian location, we draw from point cloud compression and adopt octree-based lossless compression into our framework.
-For the remaining important SHs and other attributes representing the Gaussian shape, rotation, and opacity, we save these features in float16 format.
-Vector Quantization
-VQ partitions the Gaussians $\mathcal{G}=\{\mathbf{g}_{1},\mathbf{g}_{2},\ldots,\mathbf{g}_{N}\}$ to the K codes in the codebook $\mathcal{C}=\{\mathbf{c}_{1},\mathbf{c}_{2},\ldots,\mathbf{c}_{K}\}$, where each $\mathbf{g}_{j},\mathbf{c}_{k}\in\mathbb{R}^{d}$ and K $\ll$ N.
-$d$ means attributes’ dimension.
-Specifically, $\mathcal{C}$ is initialized via K-means.
-Then, VQ iteratively samples a batch of $\mathcal{G}$, associates them to the closest codes by euclidean distance, and updates each $\mathbf{c}_{k}$ via moving average rule: $\mathbf{c}_{k}=0.8\cdot\mathbf{c}_{k}+(1-0.8)\cdot 1/\mathbf{T}_{k}\cdot\sum_{\mathbf{g}_{j}\in\mathcal{R}(\mathbf{c}_{k})}\mathbf{GS}_{j}\cdot\mathbf{g}_{j}$, where $\mathbf{T}_{k}=\sum_{\mathbf{g}_{j}\in\mathcal{R}(\mathbf{c}_{k})}{\mathbf{GS}_{j}}$ is the total significance score assigned to the code vector $\mathbf{c}_{k}$, $\mathcal{R}(\mathbf{c}_{k})$ is the set of Gaussians associated to the $k$-th code.
-Methods Mip-NeRF 360 Datasets Tank and Temple Datasets
-[…]
-Plenoxels [67] 6.79 2100MB 23.62 0.670 0.443 11.2 2700MB 21.07 0.719 0.379
-INGP-Big [45] 9.43 48MB 26.75 0.751 0.299 2.79 48MB 21.92 0.744 0.305
-Mip-NeRF 360 [5] 0.06 8.6MB 29.23 0.844 0.207 0.09 8.6MB 22.21 0.759 0.257
-VQ-DVGO [38] 6.53 56.07MB 25.44 0.699 0.325 - - - - -
-3D-GS [32] 139 727MB 29.13 0.870 0.185 106 380MB 23.11 0.822 0.219
-Ours 215 42.48MB 28.45 0.857 0.210 209 22.43MB 22.83 0.807 0.242
+3.4 Gaussian Attributes Vector Quantization
+Vector quantization clusters the voxels into compact codebooks.
+While high compression rates are attainable, the task of quantizing Gaussian attributes in a point-based representation, inherently non-Euclidean, presents notable challenges.
+Our empirical investigations reveal that quantization applied to all elements, particularly attributes such as position, rotation, and scale, tends to incur significant accuracy losses when represented discretely, leading to a marked reduction in precision.
+Vector Quantization.
+We propose applying Vector Quantization (VQ) to the Spherical Harmonics (SH) coefficients, based on the assumption that a subgroup of 3d Gaussians typically exhibits a similar appearance.
+Fundamentally, VQ segments the Gaussians $\mathcal{G}=\{\mathbf{g}_{1},\mathbf{g}_{2},\ldots,\mathbf{g}_{N}\}$ (here we apply it on SH) to the K codes in the codebook $\mathcal{C}=\{\mathbf{c}_{1},\mathbf{c}_{2},\ldots,\mathbf{c}_{K}\}$, where each $\mathbf{g}_{j},\mathbf{c}_{k}\in\mathbb{R}^{d}$ and K $\ll$ N.
+$d$ means SH dimension.
+We aim to strike a balance between rendering quality loss and compression rate by leveraging the pre-computed significance score from Eq.
+3.
+Based on this, we apply Vector Quantization (VQ) selectively on the least significant elements in the Spherical Harmonics (SHs).
+Specifically, we initialize $\mathcal{C}$ via K-means, iteratively sample a batch of $\mathcal{G}$, associates them to the closest codes by euclidean distance, and update each $\mathbf{c}_{k}$ via moving average rule: $\mathbf{c}_{k}=\lambda_{d}\cdot\mathbf{c}_{k}+(1-\lambda_{d})\cdot 1/\mathbf{T}_{k}\cdot\sum_{\mathbf{g}_{j}\in\mathcal{R}(\mathbf{c}_{k})}\mathbf{GS}_{j}\cdot\mathbf{g}_{j}$, where $\mathbf{T}_{k}=\sum_{\mathbf{g}_{j}\in\mathcal{R}(\mathbf{c}_{k})}{\mathbf{GS}_{j}}$ is the significance score (Eq.
+3) which is assigned to the code vector $\mathbf{c}_{k}$, $\mathcal{R}(\mathbf{c}_{k})$ is the set of Gaussians associated to the $k$-th code.
+$\lambda_{d}=0.8$ represents the decay value, which is utilized to update the code vector using a moving average.
+We fine-tune the codebook for 5,000 iterations, while fixing the gaussian-to-codebook mapping.
+We disable additional clone/split operations and leverage photometric loss on the training views.
+To maintain the integrity of crucial attributes such as these Spherical Harmonics with higher global significance scores, and Gaussian’s position, shape, rotation, and opacity attributes, we opt to bypass the VQ process and instead store these features directly in the float16 format.
 Table 1: Quantitative Comparisons in Real-world Large-scale Scenes.
@@
 Our method achieves a good balance among FPS, model size, and rendering quality.
+We re-train 3D-GS using their original code and configurations on our platform, to perform a fair comparison, noted as 3D-GS.
+Methods Mip-NeRF 360 Datasets Tank and Temple Datasets
+[…]
+Plenoxels [67] 6.79 2.1 GB 23.08 0.626 0.463 11.2 2.3 GB 21.08 0.719 0.379
+INGP-Big [46] 9.43 48MB 25.59 0.699 0.331 14.4 48MB 21.92 0.745 0.305
+Mip-NeRF 360 [5] 0.06 8.6MB 27.69 0.792 0.237 0.14 8.6MB 22.22 0.759 0.257
+VQ-DVGO [40] 4.65 63MB 24.23 0.636 0.393 - - - - -
+3D-GS [32] 134 734MB 27.21 0.815 0.214 154 411MB 23.14 0.841 0.183
+3D-GS [32] 119 724MB 27.53 0.810 0.227 106 380MB 23.11 0.822 0.219
+Ours 209 42MB 27.28 0.805 0.243 209 22MB 23.11 0.817 0.231
 Figure 4: Visual Comparisons.
 We compare LightGaussian with the vanilla 3D-GS [32], presenting a residual map between prediction and ground-truth scaled from 0 to 127 to highlight the differences.
-We observe that LightGaussian preserves the specular reflections (yellow boxes) after converting to a compact format.
+We observe that LightGaussian could preserve most of the specular reflection (yellow boxes) after converting to a compact format.
 Additionally, a slight lightness change is noted after the conversion, as shown in the bottom white box.
-For dynamic viewpoint comparisons, please refer to our supplementary video material.
-Octree-based Compression
-To deal with the unstructured and high-precision 3D points location, we avoid the lossy compression which is sensitive to the subsequent rasterization accuracy.
-Drawing inspiration from the Point Cloud Compression literature [52], which applies lossless encoding on the Gaussian location, utilizing the octree geometry codec in G-PCC [44] consumes only a very small amount of bits.
+Note that the residual maps plotted illustrate the differences between the rendered images and the GT RGB values.
+Here, darker regions indicate closer alignment.
+For a comprehensive dynamic viewpoint comparison, please refer to our supplementary video material.
 4 Experiments
 4.1 Experimental Settings
-Datasets and Metrics
-Comparisons are performed on the widely adopted scene-scale view synthesis data created by Mip-NeRF360 [6], which contains nine real-world large-scale scenes (5 unbounded outdoors, 4 indoors with detailed backgrounds).
-We utilize the published seven scenes for comparison following the train/test split in Mip-NeRF360 [6].
-Additionally, we use another large-scale unbounded dataset, the Tanks and Temples dataset [34], where we perform comparisons using the same two scenes as used in 3D-GS [32].
-In total, the adopted datasets consist of scenes with very different capture styles and cover both bounded indoor scenes with detailed backgrounds and large unbounded outdoor cases.
-The rendering quality is reported using metrics such as peak signal-to-noise ratio (PSNR), structural similarity (SSIM), and perceptual similarity via LPIPS [69].
-Compared Baselines
-We compare with methods that can model large-scale scenes, including Plenoxel [67], Mip-NeRF360 [5], and 3D-GS [32], along with representative efficient methods such as Instant-ngp [45], which uses a hash grid for storage, and VQ-DVGO [38], utilizing DVGO [55] as the basic representation and introducing voxel pruning and vector quantization to it.
+Datasets and Metrics.
+We conduct comparisons using the scene-scale view synthesis dataset provided by Mip-NeRF360 [6], which comprises nine real-world large-scale scenes, including five unbounded outdoor and four indoor settings with intricate backgrounds.
+Furthermore, we utilize the Tanks and Temples dataset [34], a comprehensive unbounded dataset, for additional comparisons, selecting the same scenes as those used in 3D-GS [32].
+Performance metrics on synthetic object-level datasets will be detailed in the supplementary materials due to constraints on space.
+We report metrics including the peak signal-to-noise ratio (PSNR), structural similarity (SSIM), and perceptual similarity as measured by LPIPS [70].
+Compared Baselines.
+We conduct comparisons with methodologies capable of modeling large-scale scenes.
+This includes Plenoxel [67], Mip-NeRF360 [5], and 3D-GS [32].
+Additionally, we compare against efficient methods like Instant-ngp [46], which leverages a hash grid for storage efficiency, and VQ-DVGO [40].
+The latter builds upon the DVGO framework [57], incorporating enhancements such as voxel pruning and vector quantization for optimized representation.
+Table 2: Ablation studies on the Gaussian Pruning & Recovery, SH Compactness, and the Vector Quantization.
+Scene: Room.
+Zero-shot Gaussian pruning leads to a degradation in rendering quality (#2), but Co-adaptation can recover most of the scene details (#3).
+Directly eliminating high-order SH negatively affects the quality (#4), while distillation with pseudo-view helps to mitigate the gap (#5, #6).
+Codebook quantization further significantly reduces the required model size and bandwidth (#7).
 […]
@@
 [6]     + Pseudo-views 300.60 77MB 31.59 0.923 0.211
-[7]  + Codebook Quant. 300.60 23MB 31.26 0.917 0.220
-[8]  + Octree Compression. 300.60 20MB 31.26 0.917 0.220
-[9] LightGaussian (Ours) 300.60 20MB 31.26 0.917 0.220
-Table 2: Ablation studies on the Gaussian Pruning, SH Compactness, and the VecTree Compression.
-Scene: Room.
-Zero-shot Gaussian pruning leads to a degradation in rendering quality (#2), but Co-adaptation can recover most of the scene details (#3).
-Directly eliminating high-order SH negatively affects the quality (#4), while distillation with pseudo-view helps to mitigate the gap (#5, #6).
-Codebook quantization significantly reduces the required model size (#7), while lossless Octree compression helps to preserve the overall quality.
+[7]    + Codebook Quant. 300.60 20MB 31.16 0.915 0.218
+[8]    + VQ finetune 300.60 20MB 31.67 0.923 0.212
+[9] LightGaussian (Ours) 300.60 20MB 31.67 0.923 0.212
+Table 3: Ablation study of the Gaussian Pruning & Recovery, by using different Gaussian attributes for computing its global significance score.
+By considering only the hit count of each Gaussian from training rays, the zero-shot pruning leads to inferior performance.
+Incorporating the opacity and volume drives us to a better criterion.
+The subsequent Gaussian Co-adaptation is used to recover most of the information loss from the pruning of redundant Gaussians.
 […]
@@
     + Co-adaptation 303.99 116MB 31.85 0.925 0.206
-Table 3: Ablation study of the Gaussian Pruning, by using different Gaussian attributes for computing its global significance score.
-By considering only the hit count of each Gaussian from training rays, the zero-shot pruning leads to inferior performance.
-Incorporating the opacity and volume drives us to a better criterion.
-The subsequent Gaussian Co-adaptation is used to recover most of the information loss from the pruning of redundant Gaussians.
+Table 4: Ablation study Gaussian Attribute Vector Quantization (VQ).
+By quantizing all attributes to FP16, a smaller model is achieved.
+VQ is applied to all attributes leads to inferior modeling accuracy, but this can be mitigated by using Global Significance (GS) on the least crucial Gaussians.
+Other attributes (e.g., scale) are also sensitive to VQ, hence we only apply VQ on the SH.
+Combining VQ and significance score as indicators, LightGaussian demonstrates a good balance between size and quality.
 […]
 Baseline 77.09MB 31.59 0.923 0.211
-+FP16 36.76MB 31.58 0.923 0.212
- + VQ All att. 18.58MB 22.97 0.750 0.378
- + VQ All att. $\times$ GS 18.58MB 26.39 0.830 0.327
- + VQ SH. 22.53MB 30.94 0.910 0.225
- + VQ SH $\times$ GS 22.53MB 31.26 0.917 0.220
-VecTree 20.65MB 31.26 0.917 0.220
-Table 4: Ablation study of hybrid VecTree Quantization (VQ).
-By quantizing all attributes to FP16, except for Gaussian locations due to their sensitivity, a smaller model is achieved.
-VQ is applied to all attributes leads to inferior modeling accuracy, but this can be mitigated by using Global Significance (GS) on the least crucial Gaussians.
-Other attributes (e.g., scale) are also sensitive to VQ, hence we only apply VQ on the SH.
-Combining this with Octree compression results in our VecTree Compression (the last row), which demonstrates a good balance between size and quality.
-Implementation Details
-The framework is developed using Pytorch and reuses the differential Gaussian rasterization presented by 3D-GS [32].
-We use our own trained checkpoints that match performance with those detailed in the 3D-GS paper [32].
-For the Global Significance Calculation, we set the power to 0.1 in Eq.
-5, and fine-tune for 5,000 steps in Gaussian Co-adaptation.
-For the SHs Distillation, we reduce the 3-degree SHs to 2-degree, eliminating 21 elements for each Gaussian, and enhance it with $\sigma$ set to 0.1 in the pseudo view synthesis.
-For the VecTree Compression, we set the codebook size to 8192.
++FP16 33.56MB 31.47 0.921 0.212
+   + VQ All att. 18.74MB 23.02 0.748 0.376
+   + VQ All att. $\times$ GS 18.74MB 26.33 0.831 0.325
+   + VQ SH. 19.89MB 30.85 0.910 0.222
+   + VQ SH $\times$ GS 19.89MB 31.16 0.915 0.218
+LightGaussian (w/ VQ Finetune) 19.89MB 31.67 0.923 0.212
+Implementation Details.
+Our framework is implemented using Pytorch and incorporates the differential Gaussian rasterization technique from 3D-GS [32].
+We utilize our own trained models, ensuring they align with the performance benchmarks established in the 3D-GS study [32].
+All performance are evaluated on a A6000 GPU, the same as 3D-GS.
+In the Global Significance Calculation phase, we assign a power value of 0.1 in Eq.
+4 and proceed to fine-tune the model for 5,000 steps during the Gaussian Co-adaptation process.
+For SH Distillation, we downscale the 3-degree SHs to 2-degree, thereby reducing 21 elements for each Gaussian.
+This process is further optimized by setting $\sigma$ to 0.1 in the pseudo view synthesis stage.
+In the Gaussian Vector Quantization step, the codebook size is configured to 8192, selecting Spherical Harmonics (SHs) with the least $60\%$ significance score for the vector quantization (VQ ratio), to balance the trade-off between compression efficiency and fidelity.
 Figure 5: Visual Comparisons for Ablation Study.
@@
 2.
-The final model (Exp #7, #8, #9) demonstrates close results to 3D-GS (Exp #1), while the Gaussian Co-adaptation, along with SH distillation, almost completely mitigates the information loss.
+The final model (Exp #9) demonstrates close results to 3D-GS (Exp #1), while the Gaussian Co-adaptation, along with SH distillation, almost completely mitigates the information loss.
+To highlight the difference, we visualize the SSIM between the rendered images and the GT, where a whiter output in the residual signifies closer alignment with the GT, denoting better quality, and the area with greater color intensity indicates lower quality.
 4.2 Experimental Results
-Quantitative Results
-To assess the quality and model size of different methods for novel view synthesis, we summarize the quantitative numbers in Tab.
-1, including highly efficient voxel-based NeRFs (Plenoxel[67], Instant-NGP [45]), the compact MLP-based Mip-NeRF360 [5], NeRF with vector quantization [38], and 3D Gaussian Splatting [32].
-Specifically, in the Mip-NeRF360 datasets, the NeRF-based methods, benefiting from their compact representation by using MLPs, show competitive accuracy, but the slow inference speed (0.06FPS) makes them impractical for real-world applications.
-Voxel-based NeRFs partly resolve the rendering efficiency issue, but the FPS still lags behind for practical usage, while Plenoxel requires 2.1GB for storing a large-scale scene.
-VQ-DVGO optimizes the storage issue, but their pruning and purely vector quantization make it challenging to generalize to complex large-scale scenes.
-On the other hand, 3D-GS strikes a balance in rendering quality and real-time rendering speed; however, it requires nearly gigabytes for storing a single scene.
-Our pipeline, LightGaussian, achieves the fastest rendering speed (1.55$\times$) compared to all existing methods, thanks to its pruning of insignificant Gaussians for an efficient rasterization process.
-Moreover, the compact SH representation and VecTree compression further reduce the model redundancy in 3D Gaussians, shrinking the model size from 727MB to 42.48MB on Mip-NeRF360 datasets, a 17$\times$ reduction ratio.
-LightGaussian also achieves nearly 2$\times$ faster rendering speed on the Tank & Temple datasets, while reducing the storage from 380MB to 22.43MB.
-Qualitative Results
-We compared the rendering results between 3D-GS and ours, with an emphasis on intricate details and background regions as illustrated in Fig.
-4, to highlight the differences.
-It is observable that 3D-GS and our LightGaussian achieve similar visual quality, even in thin structures within the scenes (the first and third rows).
-The Gaussian compression method slightly affects the specular effect on the surface (see the second and the fourth rows).
+Quantitative Results.
+We evaluate the performance of various methods for novel view synthesis by summarizing their quantitative metrics in Tab.
+7.
+This includes highly efficient voxel-based NeRFs such as Plenoxel[67] and Instant-NGP [46], the compact MLP-based Mip-NeRF360 [5], NeRF enhanced with vector quantization [40], and 3D Gaussian Splatting [32].
+Specifically, within the Mip-NeRF360 dataset, NeRF-based methods leverage MLPs for a compact representation, achieving competitive accuracy.
+However, their slow inference speed (0.06FPS) limits practical applications.
+Voxel-based NeRFs improve rendering efficiency but still fall short of real-time performance.
+For instance, Plenoxel demands 2.1GB for storing a large-scale scene.
+VQ-DVGO addresses storage concerns yet faces generalization issues in complex large-scale scenes due to its reliance on pruning and vector quantization alone.
+Conversely, 3D-GS offers a commendable balance between rendering quality and speed but requires substantial storage for a single scene.
+Our method, LightGaussian, surpasses existing techniques with a rendering speed exceeding 200 FPS, courtesy of its efficient rasterization process that prunes insignificant Gaussians.
+Collectively, the proposed techniques significantly reduces the redundancy in 3D Gaussian model, cutting down the size from 724MB to 42MB on the Mip-NeRF360 datasets—a reduction factor of 15$\times$.
+LightGaussian also outperforms 3D-GS in rendering speed on the Tank & Temple datasets by nearly 2$\times$ while shrinking the storage requirement from 380MB to 22MB.
+Qualitative Results.
+We conducted a comparative analysis of the rendering outcomes between 3D-GS and our LightGaussian model, focusing particularly on the intricate details and background regions, as demonstrated in Fig.
+4.
+It is evident that both 3D-GS and LightGaussian maintain comparable visual quality, even in scenes with thin structures (refer to the first and third rows).
 4.3 Ablation Studies
-We ablate components of our method by separately analyzing the proposed modules.
-Specifically, we find that Iterative Gaussian Pruning is effective in removing redundant Gaussians, SHs Distillation effectively reduces the SH’s degree, and Vector Quantization efficiently compresses the feature space, leading to a more compact and efficient representation.
-These individual contributions collectively enhance the overall performance of our framework.
-Overall Analysis
+We conducted ablation studies on our method’s components to individually assess the impact of the proposed modules.
+Our findings reveal that Gaussian Pruning and Recovery is effective in eliminating redundant Gaussians, ensuring that only significant ones contribute to the scene representation.
+SH Distillation proves to be effective in reducing the degree of Spherical Harmonics, which simplifies the lighting model without compromising too much quality.
+Additionally, Vector Quantization demonstrates its efficiency in compressing the feature space, resulting in a more streamlined and efficient representation.
+When combined, these modules significantly enhance the overall efficacy and performance of our framework.
+Overall Analysis.
 We report the performance of the investigation on the proposed modules in Tab.
@@
 We verify that our design of Gaussian Pruning & Recovery is effective in removing redundant Gaussians (Exp #1$\rightarrow$3) with negligible quality degradation while preserving rendering accuracy.
-This proves that the proposed Global Significance, based on Gaussian attributes, accurately represents the critical aspects.
-By removing the high-degree SHs and transferring the knowledge to a compact representation (Exp #3 $\rightarrow$5), our method successfully demonstrates the benefits from using soft targets and extra data from view augmentation, results in negligible changes in specular reflection In practical post-processing, the vector quantization on the least important Gaussians (Exp #7), combined with the octree-based lossless compression for its geometric position (Exp #8), showcases the advantage of adopting a hybrid compression.
+This proves that the proposed global significance, based on principle of splatting, accurately represents the critical aspects of each Gaussian.
+By removing the high-degree SHs and transferring the knowledge to a compact representation (Exp #3 $\rightarrow$5), our method successfully demonstrates the benefits from using soft targets and extra data from view augmentation, results in negligible changes in specular reflection.
+In practical post-processing, the vector quantization on the least important Gaussians (Exp #7), showcases the advantage of adopting VQ to further reduce model size.
 Figure 6: Visualization of Pruned Gaussians.
@@
 The Best Significance Criteria?
-To discover the most favorable criteria to measure the global significance of each Gaussian, we study the critical elements within each Gaussian, including its hit count with all training rays, Gaussian opacity, and the function-related volume.
-As shown in Tab.
-3, considering all these three attributes by weighting the hit count using opacity and Gaussian volume produces the most accurate rendering quality after zero-shot pruning.
-A brief Gaussian Co-adaptation (last row) could restore accuracy to its level prior to pruning.
-We visualize the rasterized images before and after pruning, as well as the pruned Gaussians, in Fig.
+To discover the most favorable criteria to measure the global significance of each Gaussian, we evaluated key characteristics within each Gaussian: the Gaussian/rays interaction count, Gaussian opacity, and the volume related to its function.
+Tab.
+3 illustrates that integrating these three factors—by weighting the hit count with opacity and Gaussian volume—yields the highest rendering quality post zero-shot pruning.
+A quick Gaussian Co-adaptation (the last row) which optimizes the remaining significant Gaussians, can effectively recover the rendering accuracy to its pre-pruning level.
+The visual comparison of rasterized images, before and after pruning, alongside the depiction of pruned Gaussians, is presented in Fig.
 6.
-SH Distillation
-Directly eliminating the high-degree component of SHs leads to an unacceptable performance drop from the full model (Exp #3 $\rightarrow$ 4), where we can also observe a significant loss of specular reflection on surfaces when changing viewpoints.
-However, by introducing knowledge distillation (Exp #4 $\rightarrow$ 5) from the full model, we can reduce its size and preserve the viewing effects.
-Additionally, adding pseudo-views during training (Exp #5 $\rightarrow$ 6) further demonstrates the effectiveness of guiding the training of a student model through the teacher model.
-VecTree Attribute Compression
-We study the effectiveness of applying quantization on Gaussian attributes, as shown in Tab.
-4.
-We find that the Gaussian location is sensitive to a low-bit representation and thus apply lossless Octree-based compression on it.
-For the rest of the attributes, we first quantize them to half-precision, and applying Vector Quantization (VQ) on all attributes leads to inferior results, but this can be mitigated by using Gaussian Global Significance.
-The final module, VecTree Compression, performs Octree-based compression on Gaussian location, VQ on the least important SHs, and FP16 on the rest, achieving the best performance balance.
-Degradation & Speed vs. Compression Ratio?
-To strike a balance among model compression ratio and the reduced model size as well as the rendering speed, we study the effect of various ratios for Gaussian Pruning & Recovery, SHs Distilling, and VecTree Compression.
-As can be seen in the Fig 7 and Tab 5.
-The ratios selected represent a balanced point to optimize the trade-off between model size, FPS, and rendering quality (see Tab 5).
-Figure 7: Visualization of The effect from different prune and VQ ratio.
+SH Distillation & Vector Quantization.
+The direct removal of high-degree components in Spherical Harmonics (SHs) results in a substantial performance decline compared to the full model (Exp #3 $\rightarrow$ 4), manifesting notably in the loss of specular reflections on surfaces with varying viewpoints.
+However, the integration of knowledge distillation (Exp #4 $\rightarrow$ 5) from the full model enables us to decrease its size while maintaining the essential viewing effects.
+Furthermore, the incorporation of pseudo-views during the training process (Exp #5 $\rightarrow$ 6) validate the efficacy of using the simulated views to enhance the learning of specular reflections.
+Applying Vector Quantization (VQ) on all attributes leads to inferior results (See Tab.
+4), but this can be mitigated by using Gaussian Global Significance.
+Table 5: Ablation under different SH ompactness: With a Gaussian pruning ratio of 66% and reduce SH to 2-degree, LightGaussian nearly preserves the rendering quality (SSIM decrease from 0.927 to 0.926).
+Reducing the SH to 1-degree lead to the SSIM decreases to 0.923.
 Exp ID Model setting Model Size $\downarrow$ FPS $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
 [1] 3D-GS 353.27 MB 192 31.687 0.927 0.200
-[2] 3D-GS (w/ 40% pruned Gaussian) 211.96 MB 245 31.930 0.928 0.198
 [2] [1] + 2-degree SH compactness 140.16 MB 238 30.625 0.917 0.208
@@
 [5] [4] + proposed distillation 88.89 MB 244 31.440 0.923 0.205
-[7] [4] + 55% vector quantization 42.92MB 244 31.626 0.924 0.204
-Table 5: Ablation under a mild compression ratio: With a Gaussian pruning ratio of 40% and VQ ratio of 55%, LightGaussian nearly preserves the rendering quality (SSIM decrease from 0.927 to 0.924), while achieving an $\sim$87reduction in model size and enhancing the FPS from 192 to 244.
-5 Conclusion
-We present LightGaussian, a novel framework that converts the heavy point-based representation into a compact format for efficient novel view synthesis.
-For practical use, LightGaussian explores the use of 3D Gaussians for modeling large-scale scenes and finds an effective way to identify the least important Gaussians grown by densification.
-In pursuing a compact Spherical Harmonics format, distillation is enhanced by synthesizing pseudo-views to generate more data for knowledge transfer.
-A VecTree compression post-processing effectively removes further redundancy by using Gaussian significance as an indicator.
-With these methods, the proposed representation reduces data redundancy by more than 15$\times$, further boosting the FPS to more than 200 FPS, with minimal rendering quality loss.
-6 Acknowledgements
-We would like to express our gratitude to Yueyu Hu from NYU for the invaluable discussion on our project.
-A7 More Technical Details
+Figure 7: Visualization of the rendering performance vs. different prune and VQ ratios. The quantize ratio is the compression ratio at vector quantization.
+Degradation & Speed vs. Compression Ratio
+We investigate the interplay between rendering quality and speed across various model compression ratios.
+Specifically, we adjust parameters such as Gaussian Pruning & Recovery, SH Distilling, and VQ.
+As depicted in Fig 7, we note a marked decline in rendering quality when the pruning ratio reaches 70%, with a more rapid deterioration as the VQ ratio approaches 65%.
+In Tab 5, we observe that 2-degree SH compactness reduces the SSIM from 0.927 to 0.926, while 1-degree SH compactness further reduces it to 0.923.
+5 More Technical Details
+Algorithm Explanation.
 We detail the procedures of LightGaussian in Algorithm 1.
@@
 3:# $\boldsymbol{G_{i}}$ with attributes XYZ, SH-3deg, Opacity, Rotation, Scale.
-4:#Gaussian Pruning and Recovery.
 […]
+5:#Gaussian Pruning and Recovery: $\mathcal{G}\mapsto\mathcal{G^{\prime}}$.
 […]
-[…]
-[…]
-9:#Distilling into Compact SHs.
-[…]
+7:$\mathcal{G^{\prime}}\leftarrow\textsc{Prune}(\mathcal{G},\mathcal{GS})$triangleright Prune Least Significant Ones of $\mathcal{G}$, based on $\mathcal{GS}$
+8:$\mathcal{G^{\prime}}\leftarrow\textsc{Recovery}(\mathcal{G^{\prime}},\mathcal{P})$$\triangleright$ Gaussian Recovery (Finetune $\mathcal{G^{\prime}}$ on $\mathcal{P}$)
+9:#Distilling into Compact SHs: Reduce from 3-degree to 2-degree
+10:$\mathcal{G^{\prime\prime}}\leftarrow\textsc{ReduceSH}(\mathcal{G^{\prime}})$triangleright Reduce the SH degree
 11:while Few Steps do $\triangleright$ SH Distillation
 12:    $\mathcal{\hat{P}}=\text{SampleView}(\mathcal{P})$triangleright Synthesize Pseudo Views
-13:    $I_{t}\leftarrow\textsc{\text{Teacher}}(\mathcal{\hat{P}})$triangleright Teacher render
-14:    $I_{s}\leftarrow\textsc{\text{Student}}(\mathcal{\hat{P}})$triangleright Student render
+13:    $I_{t}\leftarrow\textsc{\text{Teacher}}(\mathcal{\hat{P}},\mathcal{G^{\prime}})$triangleright Teacher render
+14:    $I_{s}\leftarrow\textsc{\text{Student}}(\mathcal{\hat{P}},\mathcal{G^{\prime\prime}})$triangleright Student render
 15:    $\nabla L\leftarrow\textsc{Loss}(I_{s},I_{t})$
-16:    $\mathcal{\hat{G}}\leftarrow\textsc{Adam}(\nabla L)$triangleright Backprop & Step
+16:    $\mathcal{G^{\prime\prime}}\leftarrow\textsc{Adam}(\nabla L)$triangleright Backprop & Step
 17:end while
-18:#VecTree Quantization.
-[…]
-[…]
-21:$\text{XYZ}^{\prime}\leftarrow$ OctreeCompression(XYZ)
-22:#Save Model.
-23:Save optimized model $\mathcal{\hat{G}}$ to disk.
+18:#Vector Quantization.
+19:$\mathcal{G^{\prime\prime\prime}}\leftarrow$ VQ($\mathcal{G^{\prime\prime}},\mathcal{GS}$) $\triangleright$ VQ based on Significance Score
+20:$\mathcal{G^{\prime\prime\prime}}\leftarrow$ VQFineTune($\mathcal{G^{\prime\prime\prime}}$) $\triangleright$ VQ Finetuning
+21:#Save Model.
+22:Save optimized model $\mathcal{G^{\prime\prime\prime}}$ to disk.
 Specifically, we describe how we calculate the Global Significance Score for each trained Gaussian in Algorithm 2.
@@
 1:# $\mathcal{G}$ contains all gaussians with attributes XYZ, SH-3deg, Opacity, Rotation, Scale.
-2:# $\mathcal{P}$ Sample Camera poses from training.
+2:# $\mathcal{P}$ Sample cameras from training views.
 3:function $CalGS$($\mathcal{G},\mathcal{P}$)
-4:    $GS\leftarrow 0$triangleright Init Global Significance
-5:    for all pixels $\ i\ \ \text{in renderFunc}(\mathcal{P})$ do
-6:         $H\leftarrow$ GetHitCount($\mathcal{G},i$) $\triangleright$ Gaussian Hit count
-7:         # Eq.5 in the main draft.
-8:         $GS\leftarrow GS+H\cdot Opacity^{\intercal}\cdot$ VNorm($Scale$${}^{\intercal}$
-9:    end for
-10:    return $GS$
-11:end function
-A8 More Experiment Results
+4:    ScoreList = list()
+5:    for all Gaussian ${}_{i}\ \text{in}\ \mathcal{G}$ do
+6:         $\text{score}\leftarrow 0$triangleright Init Global Significance
+7:         $\text{count}\leftarrow 0$triangleright Init Hit Count Value
+[…]
+9:             count = GetHitCount(Gi, $r_{i}$) $\triangleright$ Gaussian Hit count
+10:             # Score weighted by opacity/volume
+[…]
+12:         end for
+13:         ScoreList.append(score)
+14:    end for
+15:    return ScoreList
+16:end function
+Post-Finetuning of Vector Quantization.
+Applying vector quantization (VQ) based on the significance score facilitates a reduction in the required codebook size compared to the full model.
+Nonetheless, this method results in a decrease in accuracy as it groups the attributes into several clusters (codes).
+We observe that the SSIM decreases from 0.923 (#6) to 0.915 (#7), as illustrated in Table 2 of the main draft.
+This observation motivates us to joint optimize all other attributes and the codes post the VQ.
+Specifically, by assigning a unique code to each Gaussian, we proceed to implement differentiable finetuning using photometric loss, which is similar to the methodology employed in 3D-GS, and optimize all the attributes in conjunction with VQ.
+This approach permits precise adjustments to the codebook and the rest attributes, improving its alignment with the training images (#8, SSIM improves from 0.915 to 0.923).
+6 Detailed Experiment Results
 In addition to realistic indoor and outdoor scenes in the Mip-NeRF360 [5] and Tank and Temple datasets [34], we further evaluate our method on the synthetic Blender dataset [45], and provide a scene-wise evaluation on all datasets, accompanied by detailed visualizations.
@@
 Similar to 3D-GS [32], we start training the model using random initialization.
-Consequently, we calculate the Global Significance of each Gaussian, work to reduce the SH redundancy, and apply the VecTree Compression (codebook size set at 8192) to the learned representation.
-Overall comparisons with previous methods are listed in Table A6, where we observe our LightGaussian markedly reduces the average storage size from 52.38MB to 7.89MB, while improving the FPS from 310 to 411 with only a slight rendering quality decrease.
-Table A6: Per-scene results on Synthetic-NeRF.
+Consequently, we calculate the Global Significance of each Gaussian, work to reduce the SH redundancy, and apply the Vector Quantization (codebook size set at 8192) to the learned representation.
+Overall comparisons with previous methods are listed in Table 6, where we observe our LightGaussian markedly reduces the average storage size from 52.38MB to 7.89MB, while improving the FPS from 310 to 411 with only a slight rendering quality decrease.
+Table 6: Per-scene results on Synthetic-NeRF.
 Method Chair Drums Ficus Hotdog Lego Materials Mic Ship Avg.
@@
 We provide extra visualizations for 3D-GS [32], LightGaussian (ours), and VQ-DVGO [40], accompanied by the corresponding residual maps from the ground truth.
-As evidenced in Figure A8 and Figure A9, LightGaussian outperforms VQ-DVGO, which utilizes NeRF as a basic representation.
+As evidenced in Figure 8 and Figure 9, LightGaussian outperforms VQ-DVGO, which utilizes NeRF as a basic representation.
 Furthermore, LightGaussian achieves a comparable rendering quality to 3D-GS [32], demonstrating the effectiveness of our proposed compact representation.
 Additional Quantitative Results on Mip-NeRF360.
-Tables A7, A8, and A9 present the comprehensive error metrics compiled for our evaluation across all real-world scenes (Mip-NeRF360 and Tank and Temple datasets).
-Our method not only compresses the average model size from 727MB to 42MB, but also consistently demonstrates comparable metrics with 3D-GS on all scenes.
+Tables 7, 8, and 9 present the comprehensive error metrics compiled for our evaluation across all real-world scenes (Mip-NeRF360 and Tank and Temple datasets).
+Our method not only compresses the average model size from 724MB to 42MB, but also consistently demonstrates comparable metrics with 3D-GS on all scenes.
 LightGaussian additionally shows better rendering quality than Plenoxel, INGP, mip-NeRF360, and VQ-DVGO.
-Figure A8: Additional Visual Comparisons on the Mip-NeRF360 Datasets.
+Figure 8: Additional Visual Comparisons on the Mip-NeRF360 Datasets.
 We present the rendering results from 3D-GS [32], LightGaussian, and VQ-DVGO [40].
 The corresponding residual maps highlighting the differences between the rendered images and ground truth (GT) images are also displayed.
-Figure A9: Additional Visual Comparisons on the Mip-NeRF360 Datasets.
+Figure 9: Additional Visual Comparisons on the Mip-NeRF360 Datasets.
 We present the rendering results from 3D-GS [32], LightGaussian, and VQ-DVGO [40].
 The corresponding residual maps highlighting the differences between the rendered images and ground truth (GT) images are also displayed.
-Table A7: Quantitative Comparison (PSNR) for Mip-NeRF360 and Tank & Temple Scenes.
+Table 7: Quantitative Comparison (PSNR) on Mip-NeRF360 and Tank & Temple Scenes. We re-train 3D-GS using their original code and configurations on our platform, to perform a fair comparison.
 Methods PSNR
-bicycle garden stump room counter kitchen bonsai truck(T&T) train(T&T)
-Plenoxels 21.912 23.4947 20.661 27.594 23.624 23.420 24.669 23.221 18.927
-INGP-Big 22.171 25.069 23.466 29.690 26.691 29.479 30.685 23.383 20.456
-mip-NeRF360 24.305 26.875 26.175 31.467 29.447 31.989 33.397 24.912 19.523
-VQ-DVGO 22.089 24.119 23.455 28.423 26.084 25.930 27.985 - -
-3D-GS 25.122 27.294 26.783 31.687 29.114 31.618 32.307 24.778 21.449
-Ours 24.960 26.735 26.701 31.271 28.113 30.402 31.014 24.561 21.095
+bicycle flowers garden stump room treehill counter kitchen bonsai truck(T&T) train(T&T)
+Plenoxels 21.912 20.097 23.4947 20.661 22.248 27.594 23.624 23.420 24.669 23.221 18.927
+INGP-Big 22.171 20.652 25.069 23.466 22.373 29.690 26.691 29.479 30.685 23.383 20.456
+mip-NeRF360 24.305 21.649 26.875 26.175 22.929 31.467 29.447 31.989 33.397 24.912 19.523
+VQ-DVGO 22.092 18.934 24.127 23.428 28.502 21.440 26.035 25.459 27.990 - -
+3D-GS 25.122 21.324 27.294 26.783 31.687 22.500 29.114 31.618 32.307 24.778 21.449
+Ours 25.093 21.252 27.043 26.862 31.680 22.658 28.464 30.974 31.459 24.851 21.407
 Implementation Details of VQ-DVGO.
@@
 All other parameters are aligned with those specified in the original VQ-DVGO paper [40], ensuring a faithful replication of established methodologies.
-Table A8: Quantitative Comparison (SSIM) for Mip-NeRF360 and Tank & Temple Scenes.
+Table 8: Quantitative Comparison (SSIM) on Mip-NeRF360 and Tank & Temple Scenes. We re-train 3D-GS using their original code and configurations on our platform, to perform a fair comparison.
 Methods SSIM
-bicycle garden stump room counter kitchen bonsai truck(T&T) train(T&T)
-Plenoxels 0.496 0.6063 0.523 0.8417 0.759 0.648 0.814 0.774 0.663
-INGP-Big 0.512 0.701 0.594 0.871 0.817 0.858 0.906 0.800 0.689
-mip-NeRF360 0.685 0.809 0.745 0.910 0.892 0.917 0.938 0.857 0.660
-VQ-DVGO 0.473 0.613 0.564 0.860 0.777 0.763 0.842 - -
-3D-GS 0.746 0.856 0.770 0.914 0.914 0.932 0.946 0.863 0.781
-Ours 0.738 0.836 0.768 0.926 0.893 0.914 0.933 0.855 0.760
-Table A9: Quantitative Comparison (LPIPS) for Mip-NeRF360 and Tank & Temple Scenes.
+bicycle flowers garden stump room treehill counter kitchen bonsai truck(T&T) train(T&T)
+Plenoxels 0.496 0.431 0.606 0.523 0.509 0.842 0.759 0.648 0.814 0.774 0.663
+INGP-Big 0.512 0.486 0.701 0.594 0.542 0.871 0.817 0.858 0.906 0.800 0.689
+mip-NeRF360 0.685 0.584 0.809 0.631 0.745 0.910 0.892 0.917 0.938 0.857 0.660
+VQ-DVGO 0.473 0.355 0.614 0.563 0.861 0.476 0.777 0.760 0.842 - -
+3D-GS 0.746 0.574 0.856 0.770 0.926 0.630 0.914 0.932 0.946 0.863 0.781
+Ours 0.745 0.566 0.845 0.775 0.923 0.627 0.900 0.921 0.938 0.862 0.772
+Table 9: Quantitative Comparison (LPIPS) on Mip-NeRF360 and Tank & Temple Scenes. We re-train 3D-GS using their original code and configurations on our platform, to perform a fair comparison.
 Methods LPIPS
-bicycle garden stump room counter kitchen bonsai truck(T&T) train(T&T)
-Plenoxels 0.506 0.3864 0.503 0.4186 0.441 0.447 0.398 0.335 0.422
-INGP-Big 0.446 0.257 0.421 0.261 0.306 0.195 0.205 0.249 0.360
-mip-NeRF360 0.305 0.171 0.265 0.213 0.207 0.128 0.179 0.159 0.354
-VQ-DVGO 0.572 0.404 0.424 0.219 0.244 0.219 0.191 - -
-3D-GS 0.245 0.122 0.242 0.200 0.186 0.118 0.184 0.175 0.262
-Ours 0.265 0.155 0.261 0.220 0.218 0.147 0.204 0.188 0.296
+bicycle flowers garden stump room treehill counter kitchen bonsai truck(T&T) train(T&T)
+Plenoxels 0.506 0.521 0.3864 0.503 0.540 0.4186 0.441 0.447 0.398 0.335 0.422
+INGP-Big 0.446 0.441 0.257 0.421 0.450 0.261 0.306 0.195 0.205 0.249 0.360
+mip-NeRF360 0.305 0.346 0.171 0.265 0.347 0.213 0.207 0.128 0.179 0.159 0.354
+VQ-DVGO 0.571 0.628 0.402 0.425 0.216 0.640 0.244 0.222 0.191 - -
+3D-GS 0.245 0.381 0.122 0.242 0.200 0.366 0.186 0.118 0.184 0.175 0.262
+Ours 0.257 0.392 0.143 0.251 0.212 0.378 0.211 0.140 0.199 0.179 0.283
+7 Conclusion and Future Works
+We present LightGaussian, a novel framework that converts the heavy point-based representation into a compact format for efficient novel view synthesis.
+For practical use, LightGaussian explores the use of 3D Gaussians for modeling large-scale scenes and finds an effective way to identify the least important Gaussians grown by densification.
+In pursuing a compact Spherical Harmonics format, distillation is enhanced by synthesizing pseudo-views to generate more data for knowledge transfer.
+A 3D Gaussian Vector Quantization post-processing effectively removes further redundancy by using Gaussian significance as an indicator.
+With these methods, the proposed representation reduces data redundancy by more than 15$\times$, further boosting the FPS to more than 200 FPS, with minimal rendering quality loss.
+Exploring the possibility of zero-shot compression across various 3D-GS-based frameworks is a valuable direction for future research.

### v6

@@
 Abstract
-Recent advancements in real-time neural rendering using point-based techniques have paved the way for the widespread adoption of 3D representations.
-However, foundational approaches like 3D Gaussian Splatting come with a substantial storage overhead caused by growing the SfM points to millions, often demanding gigabyte-level disk space for a single unbounded scene, posing significant scalability challenges and hindering the splatting efficiency.
-To address this challenge, we introduce LightGaussian, a novel method designed to transform 3D Gaussians into a more efficient and compact format.
-Drawing inspiration from the concept of Network Pruning, LightGaussian identifies Gaussians that are insignificant in contributing to the scene reconstruction and adopts a pruning and recovery process, effectively reducing redundancy in Gaussian counts while preserving visual effects.
-Additionally, LightGaussian employs knowledge distillation and pseudo-view augmentation to transfer spherical harmonics efficients to a lower degree, allowing knowledge convert to more compact representations while maintaining scene appearance.
-Furthermore, we propose a Gaussian Vector Quantization based on Gaussian global significance, to quantize all redundant attributes, resulting in lower bitwidth representations with minimal accuracy losses.
-In summary, LightGaussian achieves an averaged compression rate over 15$\times$ while boosting the FPS from 119 to 209, enabling an efficient representation of complex scenes on Mip-NeRF 360, Tank & Temple datasets.
-Figure 1: Compressibility and Rendering Speed.
-We present LightGaussian to transform 3D Gaussians into a more compact representation.
-LightGaussian effectively prunes redundant Gaussians while preserving visual fidelity (on the left).
-It reduces the average storage from 724MB to 42MB and improves the FPS from 119 to 209.
+Recent advances in real-time neural rendering using point-based techniques have enabled broader adoption of 3D representations.
+However, foundational approaches like 3D Gaussian Splatting impose substantial storage overhead, as Structure-from-Motion (SfM) points can grow to millions, often requiring gigabyte-level disk space for a single unbounded scene.
+This growth presents scalability challenges and hinders splatting efficiency.
+To address this, we introduce LightGaussian, a method for transforming 3D Gaussians into a more compact format.
+Inspired by Network Pruning, LightGaussian identifies Gaussians with minimal global significance on scene reconstruction, and applies a pruning and recovery process to reduce redundancy while preserving visual quality.
+Knowledge distillation and pseudo-view augmentation then transfer spherical harmonic coefficients to a lower degree, yielding compact representations.
+Gaussian Vector Quantization, based on each Gaussian’s global significance, further lowers bitwidth with minimal accuracy loss.
+LightGaussian achieves an average 15$\times$ compression rate while boosting FPS from 144 to 237 within the 3D-GS framework, enabling efficient complex scene representation on the Mip-NeRF 360 and Tank & Temple datasets.
+The proposed Gaussian pruning approach is also adaptable to other 3D representations (e.g., Scaffold-GS), demonstrating strong generalization capabilities.
+   Project Website: https://lightgaussian.github.io
+Figure 1: Compressibility and Speed: LightGaussian compacts 3D Gaussians, reducing storage from 782MB to 45MB and boosting FPS from 144 to 237, while maintaining visual quality.
 1 Introduction
 Novel view synthesis (NVS) aims to generate photo-realistic images of a 3D scene from unobserved viewpoints, given a set of calibrated multi-view images.
-NVS holds importance as it can be used for a wide range of real-world applications, including virtual reality [12], augmented reality [71], digital twin [13], and autonomous driving [64].
-Neural Radiance Fields (NeRFs)[45, 4, 5] have demonstrated promising ability for photo-realistic 3D modeling and synthesis from multi-view images where 3D location and view directions are mapped to view-dependent color and volumetric density.
-The pixel intensity can be rendered using the volume rendering technique[14].
-However, NeRF and its variants have struggled with rendering speeds for practical deployment in real-world scenarios.
-Subsequent efforts along this trend introduce voxel-based implicit representation [57, 62, 54, 67, 10], hash grid [46], computational parallelism [51], or neural light field representation [61] to speed up the rendering.
-However, they either require specific design or face a trade-off between rendering quality and efficiency, making them difficult to generalize to large-scale scenarios with practical speed.
-Recent progress in point-based 3D Gaussian Splatting (3D-GS) [32] has brought photo-realistic rendering quality to the real-time level, even for complex scenes.
-The authors utilize the idea of representing the scene as explicit 3D Gaussians with specific properties to model the scene; the 2D images are efficiently rendered using a technique named splatting[36].
-The optimal balance between speed and quality indicates a potential trend to use 3D-GS as a new representation for generating numerous large-scale scenes in digital twins and autonomous driving.
-However, using point-based representations inevitably introduces significant storage costs, as each point and its attributes are stored independently.
-For instance, 3D-GS on a typical unbounded 360 scene [5] requires more than one gigabyte, which inhibits the scalability of 3D-GS (e.g., 1.4GB on scene Bicycle).
-In this paper, we counteract the heavy storage issue and deliver a compact representation while preserving the rendering quality.
-By examining the well-trained point-based representation, each scene is composed of millions of Gaussians, grown from sparse Structure from Motion (SfM) point cloud.
-Attributes are attached to Gaussians to model the scene’s geometry and appearance.
-However, the significant number of Gaussians after its optimization, as well as the high-degree Spherical Harmonics (SH) coefficients used for modeling scene reflection, contribute to an over-parametrized representation when fitting to the scenes.
-To minimize the required Gaussian number, we propose proper criteria to measure the global significance of each 3D Gaussian, in the context of its contribution for view synthesis.
-Gaussians that have minimal impact on visual quality are identified and pruned.
-This is followed by streamlined recovery steps that allow for rapid optimization, named as Gaussian Pruning and Recovery.
-Spherical harmonics (SH) coefficients, constituting a significant portion of the data, are employed to model view-dependent color and to capture the interaction of light with surfaces from diverse angles.
-However, utilizing higher-degree SH coefficients presents a trade-off between accuracy and computational efficiency.
-While higher-degree coefficients offer more detailed lighting information, enhancing scene realism, they also escalate computational complexity and storage demands.
-To mitigate this, we propose a SH Distillation step complemented by pseudo-view augmentation.
-This approach is designed to efficiently condense the information from higher-degree coefficients into a more compact form, ensuring that quality is preserved without inflating computational requirements.
-Besides focusing on reducing redundancy in spatial (Gaussian count) and lighting (SHs) representations, we additionally introduce a Vector Quantization step.
-This process adaptively selects a specific quantity of distinct point attributes, determined by their global significance, which further reduces the required bitwidth from the original format.
-In summary, our proposed framework, LightGaussian, efficiently reduces the Gaussian count (e.g., from 1.49M to 575K, Fig.
-1, left), significantly reducing storage requirements from 724MB to 42MB, with minimal rendering quality decrease ($\downarrow$ 0.005 in SSIM) on the Mip-NeRF 360 dataset (Fig.
-1, right).
-LightGaussian further improves the rendering speed to a higher level (200+ FPS) on complex scenes containing detailed backgrounds, suggesting a viable solution for broadening the application scope.
+This capability has widespread applications in virtual reality [1], augmented reality [2], digital twins [3], and autonomous driving [4].
+Neural Radiance Fields (NeRFs) [5, 6, 7] have shown promise in 3D modeling from multi-view images by mapping 3D locations and view directions to view-dependent color and volumetric density, with pixel intensity rendered through volume rendering [8].
+However, NeRF and its variants face limitations in rendering speed, limiting their deployment in real-world scenarios.
+To address this, voxel-based representations [9, 10, 11, 12, 13], hash grids [14], and neural light fields [15] have been developed.
+Despite improvements, these methods often compromise between quality and speed.
+Recent progress in point-based 3D Gaussian Splatting (3D-GS) [16] has enabled real-time rendering with photo-realistic quality for complex scenes.
+By representing the scene with explicit 3D Gaussians and using a splatting technique [17], 3D-GS balances speed and quality, making it suitable for large-scale scenarios like digital twins and autonomous driving.
+However, point-based methods incur high storage costs, as each point and its attributes require independent storage.
+Additionally, heuristic densification of sparse SfM points into dense Gaussians often results in overparameterization, leading to excessive storage and slower rendering speeds.
+For instance, a typical unbounded 360-degree scene in 3D-GS [7] may require over 1GB of storage (e.g., 1.4GB for the Bicycle scene).
+In this paper, we address storage and rendering speed issues by developing a compact representation that retains the original rendering quality.
+The heuristic densification process in 3D-GS results in significant redundancy.
+Our method, LightGaussian, reduces redundancy by targeting both Gaussian count (N) and feature dimension (F) through a comprehensive pipeline:
+  • To reduce Gaussian count (N), we propose a Gaussian Pruning and Recovery step, identifying and removing Gaussians with minimal impact on visual quality, followed by recovery to ensure smooth adaptation.
+  • For compressing features (F), we introduce an SH Distillation process to compact higher-degree spherical harmonic (SH) coefficients, supported by pseudo-view augmentation.
+Additionally, we employ Vector Quantization (VQ) to adaptively select a codebook of Gaussian attributes (e.g, positions, scales, and rotations), reducing precision for less significant features and applying quantization-aware fine-tuning to maintain quality.
+In summary, LightGaussian achieves substantial compression (e.g., from 782MB to 45MB) while maintaining visual fidelity (SSIM decrease of only 0.007 on Mip-NeRF 360).
+Rendering speed also improves, reaching over 200 FPS on complex scenes.
+Our Gaussian Pruning and Recovery method generalizes well, enhancing performance across different 3D Gaussian formats, such as Scaffold-GS.
 2 Related Works
-Efficient 3D Scene Representations for NVS.
-Neural radiance field
-(NeRF) [45] propose to use multi-layer perceptron (MLPs) to represent a scene, and this compact representation has brought view synthesis quality to a new stage.
-However, NeRF suffers from extremely slow inference challenges.
-Follow-ups either explore ray re-parameterizations[4, 5], explicit spatial data structures [68, 21, 41, 31, 57, 46, 10], caching and distillation [61, 22, 28, 51], or ray-based representations [2, 56] for speeding up.
-Still, NeRF-based methods struggle to achieve real-time rendering speed in practical large-scale scenes, which is caused by the multiple queries needed for rendering a single pixel, limiting their practical use.
-Point-based representations (e.g., Gaussian) have been extensively explored in various domains, including shape reconstruction [33], molecular structures modeling [7], and the replacement of point-cloud data [15].
-Additionally, these representations are utilized in shadow [49] and cloud rendering applications [43].
-Recently, Pulsar [37] has shown its efficient sphere rasterization, 3D Gaussian Splatting (3D-GS) [32] utilizes anisotropic Gaussians [72] and introduces a tile-based sorting for rendering, achieves real-time speed with comparable rendering quality to the MLP-based ones(e.g., Mip-NeRF 360 [5]).
-Although promising, the heavy storage requirement to store all attributes attached to the Gaussians often requires gigabyte-level disk space for saving a single unbounded scene, and millions of Gaussians hinder the rendering efficiency of 3D-GS.
-Concurrent works adopt region-based vector quantization [48], K-means based codebooks [47], or learned binary masks for each Gaussian, replacing the SHs with grid-based neural networks [39] within a modified training framework to reduce model size.
-Model Pruning and Vector Quantization.
-Model pruning involves reducing the complexity of a neural network by eliminating non-significant parameters to maintain a balance between performance and resource utilization.
-Unstructured [38] and structured pruning [27, 1] remove components at weight-level and neuron (mostly channel) levels to provide a smaller network with a smaller or more efficient network architecture.
-The iterative magnitude pruning (IMP) method, where weights of the smallest magnitude are progressively pruned over multiple iterations, has been highly successful in lottery ticket rewinding [19, 20].
-Additionally, vector quantization [59] aims to represent data with discrete entries of a learned codebook (i.e., tokens) to achieve lossy compression.
-In general, the mean square error (MSE) is used to find the most similar pattern in the codebook to replace the original input data vector.
-Previous works [11, 23, 44] have shown that learning a discrete and compact representation not only contributes to visual understanding but also improves the robustness of models.
-In this vein, vector quantization has been widely adopted in image synthesis [16], text-to-image generation [24], and novel view synthesis [68, 40, 25].
+Efficient 3D Scene Representations.
+Neural radiance fields (NeRF) [5] use multi-layer perceptrons (MLPs) to represent scenes, setting new standards for view synthesis quality.
+However, NeRF models face challenges with slow inference speeds, limiting practical use.
+Efforts to address this have explored ray re-parameterizations [6, 7], explicit spatial data structures [18, 19, 20, 21, 9, 14, 13], caching and distillation techniques [15, 22, 23, 24], and ray-based representations [25, 26].
+Nevertheless, achieving real-time rendering remains challenging for NeRF methods, especially for large-scale scenes where multiple queries per pixel hinder performance.
+Point-based representations, like Gaussians, have been explored in various applications, such as shape reconstruction [27], molecular modeling [28], and point-cloud replacement [29], as well as in shadow [30] and cloud rendering [31].
+Pulsar [32] demonstrated efficient sphere rasterization, while 3D Gaussian Splatting (3D-GS) [16] applies anisotropic Gaussians [33] with tile-based sorting to achieve real-time speed and quality comparable to MLP-based methods like Mip-NeRF 360 [7].
+Despite its strengths, 3D-GS has high storage demands due to the extensive attributes stored with each Gaussian, often requiring gigabytes per scene and hindering rendering efficiency.
+Recent concurrent works address these issues by using region-based vector quantization [34], K-means codebooks [35], view-direction exclusion [36], or learned binary masks and grid-based neural networks instead of spherical harmonics (SHs) [37] to reduce model size.
+Pruning and Quantization.
+Model pruning reduces neural network complexity by removing non-significant parameters, balancing performance and resource use.
+Unstructured [38] and structured pruning [39, 40] eliminate parameters at the weight level and neuron/channel levels, resulting in a smaller, more efficient architecture.
+Iterative magnitude pruning (IMP), where low-magnitude weights are progressively removed, has proven effective in methods like lottery ticket rewinding [41, 42].
+Vector quantization (VQ) [43] compresses data by representing it with discrete codebook entries (tokens), using mean square error (MSE) to select the closest match in the codebook for each data vector.
+Prior studies [44, 45, 46] have shown that learning discrete, compact representations enhances visual understanding and model robustness.
+VQ has thus been widely applied in image synthesis [47], text-to-image generation [48], and novel view synthesis [18, 49, 50].
 Knowledge Distillation.
-Knowledge Distillation (KD) is widely adopted in various machine learning tasks [55, 53, 63, 61].
-A student model is trained for the primary purpose of model compression [8] or ensembling [30] with the help of teacher models.
-Many variants [3, 52, 60, 69, 29, 42, 58] have been proposed to achieve better knowledge transfer from the teacher to the student models.
-In the field of 3D vision, neural scene representations have embraced knowledge distillation mainly through view renderings to leverage existing 2D priors.
-DreamFusion [50] and NeuralLift-360 [65] adopt pre-trained text-to-image diffusion models for 3D generation.
-DFF [35], NeRF-SOS [18], INS [17], SA3D [9] distill various pre-trained 2D image feature extractors to perform corresponding tasks in the 3D domain.
-Knowledge distillation has also played a key role in model compression of scene representations.
-R2L [61] and KiloNeRF [51] distill Neural Radiance Fields (NeRF) into more efficient representations such as light fields or multiple tiny MLPs.
-Our work falls into the category of utilizing knowledge distillation for model compression and leverages novel view renderings as the bridge to connect the teacher and student models.
+Knowledge Distillation (KD) [51, 52, 53, 54, 15] trains a smaller student model by transferring knowledge from a larger teacher model [55].
+In 3D vision, KD has been applied to neural scene representations, leveraging view renderings to incorporate 2D priors.
+For example, DreamFusion [56] and NeuralLift-360 [57] use pre-trained diffusion models for 3D generation, while models like DFF [58], NeRF-SOS [59], INS [60], and SA3D [61] distill 2D image feature extractors to 3D tasks.
+KD has also been central to compressing scene representation models [15, 24].
 3 Methods
-Figure 2: The overall pipeline of LightGaussian.
+Figure 2: Pipeline of LightGaussian.
 3D Gaussians are optimized from multi-view images and SfM points.
-LightGaussian first calculates the global significance for each Gaussian on training data, where Gaussians with the least significance are pruned.
-A subsequent distillation with synthesized pseudo-views is introduced to transfer the SH into a compact format.
-Vector quantization, consisting of codebook initialization and assignment, is further introduced for reducing the model bandwidth.
+LightGaussian calculates each Gaussian’s global significance on training data, pruning those with the least significance.
+Next, SH coefficients are distilled into a compact format using synthesized pseudo-views.
+Finally, vector quantization, including codebook initialization and assignment, reduces model bandwidth.
 Overview.
-The overview of LightGaussian is illustrated in Fig.
+The LightGaussian framework is illustrated in Fig.
 2.
-The 3D-GS model is trained using multi-view images and is initialized from SfM point clouds.
-By expanding the sparse points to millions of Gaussians, the scene is well-represented.
-Then, the 3D-GS undergoes processing within our pipeline to transform it into a more compact format.
-This involves utilizing Gaussian Prune and Recovery to reduce the number of Gaussians, SH Distillation to remove redundant SHs while preserving the essential lighting information, and Vector Quantization to store Gaussians at a lower bit-width.
+The 3D-GS model is trained on multi-view images, initialized from SfM point clouds, and expanded to millions of Gaussians to represent the scene comprehensively.
+Our pipeline then processes the 3D-GS model into a compact format using Gaussian Prune and Recovery to reduce the number of Gaussians, SH Distillation to eliminate redundant SHs while retaining key lighting information, and Vector Quantization to store Gaussians with lower bit-width.
 3.1 Background: 3D Gaussian Splatting
@@
 $\text{G}(\boldsymbol{X})=e^{-\frac{1}{2}\boldsymbol{X}^{T}\boldsymbol{\Sigma}^{-1}\boldsymbol{X}},\boldsymbol{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T},$ (1)
-where $\boldsymbol{\Sigma}$ can be decomposed into a scaling matrix $\mathbf{S}$ and a rotation matrix $\mathbf{R}$ for differentiable optimization.
+where $\boldsymbol{\Sigma}$ can be decomposed into a scaling matrix $\mathbf{S}$ and a rotation matrix $\mathbf{R}$.
 The complex directional appearance is modeled by an additional property, Spherical Harmonics (SH), with $n$ coefficients, $\left\{\boldsymbol{c}_{i}\in\mathbb{R}^{3}|i=1,2,\ldots,n\right\}$ where $n=D^{2}$ represents the number of coefficients of SH with degree $D$.
@@
 3.2 Gaussian Pruning & Recovery
-Gaussian densification [32], which involves cloning and splitting the initial SfM point cloud, is employed to address the challenge of insufficient coverage and is used to model small-scale geometry as well as detailed scene appearance.
-While this strategy leads to significantly improved reconstruction quality, it results in the number of Gaussians growing from thousands to millions after optimization.
-Such an explicit point-based representation with a large number of Gaussians requires an extremely significant storage overhead.
-Drawing inspiration from the success of representative neural network pruning techniques [26], which eliminates less impactful neurons without compromising the network’s overall performance, we tailor a general pruning paradigm for 3D Gaussian Splatting representation to reduce the over-parameterized point number in a manner that preserves the original accuracy.
-Therefore, identifying the most representative redundant Gaussians with recoverable accuracy is a crucial step in our approach.
-However, pruning the Gaussians based on simplistic criteria (e.g., point opacity) can lead to a substantial degradation in modeling performance, especially where the intricate scene structure may be eliminated, as is demonstrated in Fig.
+Gaussian densification [16], which clones and refines the initial SfM point cloud, enhances small-scale geometry and detailed scene appearance by improving coverage.
+While this approach significantly boosts reconstruction quality, it increases the number of Gaussians from thousands to millions after optimization, resulting in substantial storage demands.Inspired by neural network pruning techniques [63] that remove less impactful neurons while preserving overall performance, we propose a tailored pruning strategy for 3D Gaussian Splatting to reduce over-parameterized points while maintaining accuracy.
+Identifying redundant yet recoverable Gaussians is essential to our approach.
+However, pruning Gaussians based on simple criteria (e.g., point opacity) risks degrading modeling performance, as it may eliminate essential scene details, as shown in Fig.
 3.
@@
 Global Significance Calculation.
-Rather than merely depending on Gaussian opacity for assessing significance, we are inspired by Eq.
-2.
-In this context, 3D Gaussians within the view frustum are evaluated and rasterized by projecting them onto a camera viewpoint for rendering.
+Inspired by Eq.
+2, we go beyond using Gaussian opacity alone to assess significance by evaluating 3D Gaussians within the view frustum, projecting them onto the camera viewpoint for rendering.
 The initial significance score of each Gaussian (denoted as $\text{G}(\boldsymbol{X}_{j})$) can then be quantified based on its contribution to each pixel (ray, $r_{i}$) across all training views using the criteria $\mathbbm{1}(\text{G}(\boldsymbol{X}_{j}),r_{i})$ whether they intersect or not.
@@
 The score is further refined by the adjusted 3D Gaussian’s volume $\large{\gamma}$$(\boldsymbol{\Sigma}_{j})$ and 3D Gaussian’s opacity $\sigma_{j}$, as they all contribute for the rendering formula.
-The volume calculation equation of the $j$-th Guassian is V${(\boldsymbol{\Sigma}_{j})}=\frac{4}{3}\pi abc$, where $abc$ are the 3 dimensions of Scale ($\mathbf{S}$).
-Finally, the global significant score can be summarized as:
-$\text{GS}_{\text{j}}$ $=\sum_{i=1}^{MHW}\mathbbm{1}(\text{G}(\boldsymbol{X}_{j}),r_{i})\cdot\sigma_{j}\cdot{\large{\gamma}}(\boldsymbol{\Sigma}_{j}),$ (3)
-where $j$ is the Gaussian index, $i$ means a pixel, $M$, $H$, and $W$ represents the number of training views, image height, and width, respectively.
-$\mathbbm{1}$ is the indicator function that determines whether a Gaussian intersects with a given ray.
-However, the use of Gaussian volume tends to exaggerate the importance of background Gaussians, leading to the immoderate pruning of Gaussians that model intricate geometry.
-Therefore, we propose a more adaptive way to measure the dimension of its volume:
+The volume calculation equation of the $j$-th 3D Guassian is V${(\boldsymbol{\Sigma}_{j})}=\frac{4}{3}\pi abc$, where $abc$ are the 3 dimensions of Scale ($\mathbf{S}$).
+We also consider the transmittance $\mathbf{T}$ is calculated with one subtract all the 3D Gaussian’s opacity the ray hit before the $j$-th 3D Gaussian, $\mathbf{T}={\prod_{i=1}^{j-1}}$ (1- $\sigma_{i}$).
+Finally, the global significance score can be summarized as:
+$\text{GS}_{\text{j}}$ $=\sum_{i=1}^{MHW}\mathbbm{1}(\text{G}(\boldsymbol{X}_{j}),r_{i})\cdot\sigma_{j}\cdot\mathbf{T}\cdot{\large{\gamma}}(\boldsymbol{\Sigma}_{j}),$ (3)
+where $j$ is the Gaussian index, $i$ denotes a pixel, and $M$, $H$, and $W$ represent the number of training views, image height, and width, respectively.
+$\mathbbm{1}$ is an indicator function that determines whether a Gaussian intersects a given ray.
+However, using Gaussian volume alone tends to overemphasize background Gaussians, leading to excessive pruning of Gaussians modeling fine geometry.
+Thus, we propose a more adaptive approach to measuring volume dimensions:
 $\gamma(\boldsymbol{\Sigma})$ $=(\text{V}_{\text{norm}})^{\beta},$ (4)
-$\text{V}_{\text{norm}}$ $=\text{min}\left(\text{max}\left(\frac{\text{V}(\boldsymbol{\Sigma})}{\text{V}_{\text{max90}}},0\right),1\right).$
+$\text{V}_{\text{norm}}$ $=\text{min}\left(\frac{\text{V}(\boldsymbol{\Sigma})}{\text{V}_{\text{max90}}},1\right).$
 Here, the calculated Gaussian volume is firstly normalized by the 90% largest of all sorted Gaussians, clipping the range between 0 and 1, to avoid excessive floating Gaussians derived from vanilla 3D-GS.
@@
 Gaussian Co-adaptation.
-We rank all the Gaussians based on the computed global significance scores, which provides a quantitative basis for pruning Gaussians with lower scores.
-The remaining ones are optimized (adapted) to jointly adjust their attributes, without further densification, to compensate for the slight loss incurred by pruning.
-Note that, we utilize photometric loss on original training views with 5,000 iterations which is aligned with 3D-GS training.
+We rank all Gaussians by their global significance scores to quantitatively guide pruning of the lower-ranked Gaussians.
+The remaining Gaussians are then jointly adapted by fine-tuning their attributes—without additional densification—to offset the minor loss from pruning.
+This adaptation is performed using photometric loss on the original training views, aligned with 3D-GS training for 5,000 iterations.
 3.3 Distilling into Compact SHs
-In the uncompressed Gaussian Splat data, a considerable portion is occupied by Spherical Harmonics (SH) coefficients, necessitating (45+3) floating-point values per splat, which account for 81.3 percent of the total attribute volume.
-While directly diminishing the degree of SH is advantageous for reducing disk space consumption, this action results in a significant reduction in surface “shininess”, notably impacting the variance in specular reflections as the viewpoint changes.
-To strike a balance between model size and scene modeling quality, we suggest transferring knowledge from well-trained high-degree SHs to their more compact, lower-degree counterparts through a distillation process.
-Specifically, we distill the knowledge from the teacher model, which employs full-degree SHs and remains static, to the student model, characterized by truncated lower-degree SHs.
-The supervision is based on the differences in the predicted pixel intensities, denoted as $\boldsymbol{C}$, between the two models:
-$\mathcal{L}_{\text{distill}}=\frac{1}{HW}\sum_{i=1}^{HW}\left\lVert\boldsymbol{C}_{\text{teacher}}(r_{i})-\boldsymbol{C}_{\text{student}}(r_{i})\right\rVert_{2}^{2}.$ (5)
-Simply reintroducing these view-dependent visual effects to the network does not significantly enhance its knowledge.
-We, therefore, suggest employing the concept of data augmentation.
-This involves training the Gaussians not only to represent known views but also to learn behaviors from unseen (pseudo) views modeled by the teacher model, thus broadening their representational capacity.
-Synthesize Pseudo Views.
-Specular reflection rays are reflected off the surface of an object when the viewpoints are altered.
-To enrich the training data, we can introduce additional viewpoints through sampling, which are capable of capturing such reflectance characteristics.
-Given our knowledge of the camera locations and viewing directions for all training views, we strategically sample pseudo views around each training view, adhering to a Gaussian distribution, while maintaining consistent camera viewing directions:
+In the uncompressed Gaussian Splat representation, a substantial 81.3 percent of the feature dimension is occupied by Spherical Harmonics (SH) coefficients, requiring (45+3) floating-point values per splat.
+Directly reducing the SH degree can save disk space but also diminishes surface “shininess” and affects specular reflections.
+To balance model size with scene quality, we introduce a knowledge distillation approach.
+Knowledge is distilled from a teacher model with full-degree SHs to a student model with truncated, lower-degree SHs.
+Supervision is based on the difference in predicted pixel intensities between the two models, with images synthesized at camera positions by sampling around each training view according to a Gaussian distribution:
+$\mathcal{L}_{\text{distill}}$ $=\frac{1}{HW}\sum_{i=1}^{HW}\left\lVert\boldsymbol{C}_{\text{teacher}}(r_{i};[\mathbf{R}|\mathbf{t}])-\boldsymbol{C}_{\text{student}}(r_{i};[\mathbf{R}|\mathbf{t}])\right\rVert_{2}^{2}.$ (5)
 $\mathbf{t}_{\text{pseudo}}$ $=\mathbf{t}_{\text{train}}+\mathcal{N}(0,\sigma^{2}),$ (6)
-where $\mathbf{t}_{\text{pseudo}}$ and $\mathbf{t}_{\text{train}}$ represent the newly synthesized and training camera positions, respectively.
+where $\mathbf{R}$ and $\mathbf{t}$ denote rendering camera rotation and position, $\mathbf{t}_{\text{pseudo}}$ and $\mathbf{t}_{\text{train}}$ represent the newly synthesized and training camera positions, respectively.
 $\mathcal{N}$ denotes a Gaussian distribution with mean 0 and variance $\sigma^{2}$, which is added to the original position to generate the new position.
 3.4 Gaussian Attributes Vector Quantization
-Vector quantization clusters the voxels into compact codebooks.
-While high compression rates are attainable, the task of quantizing Gaussian attributes in a point-based representation, inherently non-Euclidean, presents notable challenges.
-Our empirical investigations reveal that quantization applied to all elements, particularly attributes such as position, rotation, and scale, tends to incur significant accuracy losses when represented discretely, leading to a marked reduction in precision.
-Vector Quantization.
-We propose applying Vector Quantization (VQ) to the Spherical Harmonics (SH) coefficients, based on the assumption that a subgroup of 3d Gaussians typically exhibits a similar appearance.
+Vector quantization (VQ) clusters voxels into compact codebooks, enabling high compression rates.
+However, quantizing Gaussian attributes in a point-based, inherently non-Euclidean representation poses notable challenges.
+Our empirical findings indicate that applying quantization across all elements—especially for attributes like position, rotation, and scale—results in significant accuracy losses and a marked reduction in precision when represented discretely.
+We propose applying VQ to the Spherical Harmonics (SH) coefficients, based on the assumption that a subgroup of 3d Gaussians typically exhibits a similar appearance.
 Fundamentally, VQ segments the Gaussians $\mathcal{G}=\{\mathbf{g}_{1},\mathbf{g}_{2},\ldots,\mathbf{g}_{N}\}$ (here we apply it on SH) to the K codes in the codebook $\mathcal{C}=\{\mathbf{c}_{1},\mathbf{c}_{2},\ldots,\mathbf{c}_{K}\}$, where each $\mathbf{g}_{j},\mathbf{c}_{k}\in\mathbb{R}^{d}$ and K $\ll$ N.
@@
 3.
-Based on this, we apply Vector Quantization (VQ) selectively on the least significant elements in the Spherical Harmonics (SHs).
+Based on this, we apply VQ selectively on the least significant elements in the Spherical Harmonics (SHs).
 Specifically, we initialize $\mathcal{C}$ via K-means, iteratively sample a batch of $\mathcal{G}$, associates them to the closest codes by euclidean distance, and update each $\mathbf{c}_{k}$ via moving average rule: $\mathbf{c}_{k}=\lambda_{d}\cdot\mathbf{c}_{k}+(1-\lambda_{d})\cdot 1/\mathbf{T}_{k}\cdot\sum_{\mathbf{g}_{j}\in\mathcal{R}(\mathbf{c}_{k})}\mathbf{GS}_{j}\cdot\mathbf{g}_{j}$, where $\mathbf{T}_{k}=\sum_{\mathbf{g}_{j}\in\mathcal{R}(\mathbf{c}_{k})}{\mathbf{GS}_{j}}$ is the significance score (Eq.
@@
 We disable additional clone/split operations and leverage photometric loss on the training views.
-To maintain the integrity of crucial attributes such as these Spherical Harmonics with higher global significance scores, and Gaussian’s position, shape, rotation, and opacity attributes, we opt to bypass the VQ process and instead store these features directly in the float16 format.
-Table 1: Quantitative Comparisons in Real-world Large-scale Scenes.
-We compare LightGaussian with the original 3D-GS [32], efficient voxel-based NeRFs [67, 46], voxel NeRF with vector quantization [40], and compact MLP-based NeRF [5].
-Voxel-based methods all exhibit a lack of sufficient capacity for representing large-scale scenes and are not able to run at real-time speed.
-Mip-NeRF 360 produces the best visual quality, but it requires more than 16s to render a single image.
-Our method achieves a good balance among FPS, model size, and rendering quality.
-We re-train 3D-GS using their original code and configurations on our platform, to perform a fair comparison, noted as 3D-GS.
+To preserve essential attributes, including Spherical Harmonics with higher global significance scores, along with Gaussian position, shape, rotation, and opacity, we skip VQ for these elements and store them directly in float16 format.
 Methods Mip-NeRF 360 Datasets Tank and Temple Datasets
@@
 VQ-DVGO [49] 4.65 63MB 24.23 0.636 0.393 - - - - -
+Compressed 3D-GS [34] 152 28MB 27.03 0.802 0.238 202 17MB 23.54 0.838 0.189
+Compact 3D-GS [37] 128 48MB 27.08 0.798 0.247 185 39MB 23.32 0.831 0.201
 3D-GS [16] 134 734MB 27.21 0.815 0.214 154 411MB 23.14 0.841 0.183
-3D-GS [32] 119 724MB 27.53 0.810 0.227 106 380MB 23.11 0.822 0.219
-Ours 209 42MB 27.28 0.805 0.243 209 22MB 23.11 0.817 0.231
+3D-GS [16] 144 782MB 27.40 0.813 0.217 106 433MB 23.66 0.845 0.178
+Ours 237 45MB 27.13 0.806 0.237 357 25MB 23.44 0.832 0.202
+Table 1: Quantitative Comparisons in Real-world Large-scale Scenes: Voxel-based methods [12, 14] exhibit insufficient capacity for representing large-scale scenes and are unable to achieve real-time performance.
+Mip-NeRF360 [7] produces the highest visual quality, but requires over 16 seconds to render a single image.
+Our method strikes a balance among FPS, model size, and rendering quality, achieving the best balance among efficient representations.
+For fair metric and visual comparison, we re-trained 3D-GS and Compressed 3D-GS [34] using their original code and configurations on our platform (NVIDIA A6000 GPU), with results marked by .
 Figure 4: Visual Comparisons.
-We compare LightGaussian with the vanilla 3D-GS [32], presenting a residual map between prediction and ground-truth scaled from 0 to 127 to highlight the differences.
-We observe that LightGaussian could preserve most of the specular reflection (yellow boxes) after converting to a compact format.
-Additionally, a slight lightness change is noted after the conversion, as shown in the bottom white box.
-Note that the residual maps plotted illustrate the differences between the rendered images and the GT RGB values.
-Here, darker regions indicate closer alignment.
-For a comprehensive dynamic viewpoint comparison, please refer to our supplementary video material.
+We compare LightGaussian with the vanilla 3D-GS [16], displaying a residual map between predictions and ground truth scaled from 0 to 127 to emphasize differences.
+LightGaussian retains most specular reflections (yellow boxes) in its compact format, with a slight change in lightness visible in the bottom white box.
+The residual maps illustrate discrepancies between rendered images and ground-truth RGB values, where darker areas indicate closer alignment.
+For a full dynamic viewpoint comparison, please see our supplementary video.
 4 Experiments
@@
 Datasets and Metrics.
-We conduct comparisons using the scene-scale view synthesis dataset provided by Mip-NeRF360 [6], which comprises nine real-world large-scale scenes, including five unbounded outdoor and four indoor settings with intricate backgrounds.
-Furthermore, we utilize the Tanks and Temples dataset [34], a comprehensive unbounded dataset, for additional comparisons, selecting the same scenes as those used in 3D-GS [32].
-Performance metrics on synthetic object-level datasets will be detailed in the supplementary materials due to constraints on space.
+We conduct comparisons using the scene-scale view synthesis dataset provided by Mip-NeRF360 [64], which comprises nine real-world large-scale scenes, including five unbounded outdoor and four indoor settings with complex backgrounds.
+In addition, we utilize the Tanks and Temples dataset [65], a comprehensive unbounded dataset, and select the same scenes as used in [16].
+Performance metrics on synthetic object-level datasets will be detailed in the supplementary materials.
 We report metrics including the peak signal-to-noise ratio (PSNR), structural similarity (SSIM), and perceptual similarity as measured by LPIPS [66].
 Compared Baselines.
-We conduct comparisons with methodologies capable of modeling large-scale scenes.
-This includes Plenoxel [67], Mip-NeRF360 [5], and 3D-GS [32].
-Additionally, we compare against efficient methods like Instant-ngp [46], which leverages a hash grid for storage efficiency, and VQ-DVGO [40].
-The latter builds upon the DVGO framework [57], incorporating enhancements such as voxel pruning and vector quantization for optimized representation.
+We compare our approach with methods suited for large-scale scene modeling, including Plenoxel [12], Mip-NeRF360 [7], and 3D-GS [16].
+Additionally, we evaluate against efficient techniques like Instant-NGP [14], which uses a hash grid for storage efficiency, and VQ-DVGO [49], which extends DVGO [9] with voxel pruning and VQ for optimized representation.
+[…]
+[1] Baseline (3D-GS [16]) 156.21 365MB 31.34 0.917 0.221
+[2]    + Gaussian Pruning 308.10 123MB 30.67 0.910 0.234
+[3]     + Co-adaptation 304.64 123MB 31.64 0.918 0.228
+[4]    + SH Compactness 311.58 81MB 30.32 0.904 0.238
+[5]     + photometric loss 302.20 81MB 31.42 0.916 0.234
+[6]     + Distillation + Pseudo-views 302.89 81MB 31.48 0.917 0.231
+[7]    + Codebook Quant. 301.21 21MB 31.09 0.918 0.236
+[8]    + VQ finetune 302.01 21MB 31.40 0.916 0.232
+[9] LightGaussian (Ours) 302.01 21MB 31.40 0.916 0.232
 Table 2: Ablation studies on the Gaussian Pruning & Recovery, SH Compactness, and the Vector Quantization.
 Scene: Room.
-Zero-shot Gaussian pruning leads to a degradation in rendering quality (#2), but Co-adaptation can recover most of the scene details (#3).
+Zero-shot Gaussian pruning leads degraded rendering quality (#2), but Co-adaptation can recover most of the scene details (#3).
 Directly eliminating high-order SH negatively affects the quality (#4), while distillation with pseudo-view helps to mitigate the gap (#5, #6).
-Codebook quantization further significantly reduces the required model size and bandwidth (#7).
+Codebook quantization further reduces the required model size and bandwidth (#7, #8).
 […]
-[1] Baseline (3D-GS [32]) 192.05 353MB 31.68 0.926 0.200
-[2]    + Gaussian Pruning. 312.30 116MB 30.32 0.911 0.222
-[3]     + Co-adaptation 303.99 116MB 31.85 0.925 0.206
-[4]    + SH Compactness. 318.97 77MB 30.54 0.914 0.217
-[5]     + Distillation 304.20 77MB 31.47 0.922 0.211
-[6]     + Pseudo-views 300.60 77MB 31.59 0.923 0.211
-[7]    + Codebook Quant. 300.60 20MB 31.16 0.915 0.218
-[8]    + VQ finetune 300.60 20MB 31.67 0.923 0.212
-[9] LightGaussian (Ours) 300.60 20MB 31.67 0.923 0.212
+Baseline 156.21 365MB 31.34 0.917 0.221
+Hit Count Only 301.52 123MB 28.16 0.886 0.261
+    + Co-adaptation 303.43 123MB 30.13 0.912 0.238
+   $\times$ Opacity 310.29 123MB 30.27 0.909 0.239
+    + Co-adaptation 304.84 123MB 31.60 0.916 0.231
+   $\times$ Opacity $\times$ $\gamma$(Volume). 312.30 123MB 30.67 0.910 0.234
+    + Co-adaptation 304.64 123MB 31.64 0.918 0.228
 Table 3: Ablation study of the Gaussian Pruning & Recovery, by using different Gaussian attributes for computing its global significance score.
@@
 […]
-Baseline 192.05 353MB 31.68 0.926 0.200
-Hit Count Only 300.57 116MB 28.28 0.895 0.238
-   $\times$ Opacity. 310.29 116MB 30.10 0.910 0.222
-   $\times$ Opacity $\times$ $\gamma$(Volume). 312.30 116MB 30.32 0.911 0.222
-    + Co-adaptation 303.99 116MB 31.85 0.925 0.206
-Table 4: Ablation study Gaussian Attribute Vector Quantization (VQ).
-By quantizing all attributes to FP16, a smaller model is achieved.
-VQ is applied to all attributes leads to inferior modeling accuracy, but this can be mitigated by using Global Significance (GS) on the least crucial Gaussians.
-Other attributes (e.g., scale) are also sensitive to VQ, hence we only apply VQ on the SH.
-Combining VQ and significance score as indicators, LightGaussian demonstrates a good balance between size and quality.
-[…]
-Baseline 77.09MB 31.59 0.923 0.211
-+FP16 33.56MB 31.47 0.921 0.212
-   + VQ All att. 18.74MB 23.02 0.748 0.376
-   + VQ All att. $\times$ GS 18.74MB 26.33 0.831 0.325
-   + VQ SH. 19.89MB 30.85 0.910 0.222
-   + VQ SH $\times$ GS 19.89MB 31.16 0.915 0.218
-LightGaussian (w/ VQ Finetune) 19.89MB 31.67 0.923 0.212
+Baseline 80.99MB 31.48 0.917 0.231
++FP16 36.51MB 31.35 0.914 0.232
+   + VQ All att. 18.74MB 23.11 0.731 0.378
+   + VQ All att. $\times$ GS 19.74MB 26.23 0.826 0.323
+   + VQ SH. 21.10MB 30.68 0.907 0.244
+   + VQ SH $\times$ GS 21.10MB 31.16 0.911 0.235
+LightGaussian (w/ VQ Finetune) 21.10MB 31.40 0.916 0.232
+Table 4: Ablation Study: Gaussian Attribute Vector Quantization (VQ).
+Quantizing all attributes to FP16 achieves a smaller model, but applying VQ to all attributes degrades modeling accuracy.
+Using Global Significance (GS) to target less crucial Gaussians mitigates this loss.
+Some attributes (e.g., scale) are sensitive to VQ, so we limit VQ to the SH coefficients.
+By combining VQ with significance scores, LightGaussian achieves an effective balance between model size and quality.
 Implementation Details.
-Our framework is implemented using Pytorch and incorporates the differential Gaussian rasterization technique from 3D-GS [32].
-We utilize our own trained models, ensuring they align with the performance benchmarks established in the 3D-GS study [32].
-All performance are evaluated on a A6000 GPU, the same as 3D-GS.
+Our framework is implemented in PyTorch and integrates the differentiable Gaussian rasterization technique from 3D-GS [16].
+All performance evaluations are conducted on an A6000 GPU.
 In the Global Significance Calculation phase, we assign a power value of 0.1 in Eq.
@@
 For SH distillation, we downscale the 3-degree SHs to 2-degree, thereby reducing 21 elements for each Gaussian.
-This process is further optimized by setting $\sigma$ to 0.1 in the pseudo view synthesis stage.
-In the Gaussian Vector Quantization step, the codebook size is configured to 8192, selecting Spherical Harmonics (SHs) with the least $60\%$ significance score for the vector quantization (VQ ratio), to balance the trade-off between compression efficiency and fidelity.
-Figure 5: Visual Comparisons for Ablation Study.
-We visualize the rendered RGB images and the residual map between the ground-truth image, aligned with the experiment ID as shown in Tab.
-2.
-The final model (Exp #9) demonstrates close results to 3D-GS (Exp #1), while the Gaussian Co-adaptation, along with SH distillation, almost completely mitigates the information loss.
-To highlight the difference, we visualize the SSIM between the rendered images and the GT, where a whiter output in the residual signifies closer alignment with the GT, denoting better quality, and the area with greater color intensity indicates lower quality.
+This is further optimized by setting $\sigma$ to 0.1 in the pseudo view synthesis stage.
+In the Gaussian VQ step, the codebook size is configured to 8192, selecting SHs with the least $60\%$ significance score for the vector quantization (VQ ratio), to balance the trade-off between compression efficiency and fidelity.
 4.2 Experimental Results
 Quantitative Results.
-We evaluate the performance of various methods for novel view synthesis by summarizing their quantitative metrics in Tab.
-7.
-This includes highly efficient voxel-based NeRFs such as Plenoxel[67] and Instant-NGP [46], the compact MLP-based Mip-NeRF360 [5], NeRF enhanced with vector quantization [40], and 3D Gaussian Splatting [32].
-Specifically, within the Mip-NeRF360 dataset, NeRF-based methods leverage MLPs for a compact representation, achieving competitive accuracy.
-However, their slow inference speed (0.06FPS) limits practical applications.
-Voxel-based NeRFs improve rendering efficiency but still fall short of real-time performance.
-For instance, Plenoxel demands 2.1GB for storing a large-scale scene.
-VQ-DVGO addresses storage concerns yet faces generalization issues in complex large-scale scenes due to its reliance on pruning and vector quantization alone.
-Conversely, 3D-GS offers a commendable balance between rendering quality and speed but requires substantial storage for a single scene.
-Our method, LightGaussian, surpasses existing techniques with a rendering speed exceeding 200 FPS, courtesy of its efficient rasterization process that prunes insignificant Gaussians.
-Collectively, the proposed techniques significantly reduces the redundancy in 3D Gaussian model, cutting down the size from 724MB to 42MB on the Mip-NeRF360 datasets—a reduction factor of 15$\times$.
-LightGaussian also outperforms 3D-GS in rendering speed on the Tank & Temple datasets by nearly 2$\times$ while shrinking the storage requirement from 380MB to 22MB.
+We assess the performance of various methods for novel view synthesis, with quantitative metrics summarized in Tab.
+1.
+This includes efficient voxel-based NeRFs like Plenoxel [12] and Instant-NGP [14], the compact MLP-based Mip-NeRF360 [7], vector-quantized NeRF [49], and 3D Gaussian Splatting [16].
+On the Mip-NeRF360 dataset, MLP-based NeRF methods achieve competitive accuracy with compact representations but suffer from slow inference speeds (0.06 FPS), limiting their practicality.
+Voxel-based NeRFs improve rendering efficiency but still fall short of real-time performance; for example, Plenoxel requires 2.1GB for a single large-scale scene.
+In contrast, 3D-GS offers a good balance between quality and speed but demands substantial storage per scene.
+Our method, LightGaussian, exceeds existing techniques with rendering speeds over 200 FPS, enabled by efficient rasterization that prunes insignificant Gaussians.
+This approach reduces 3D Gaussian model redundancy, cutting storage from 782MB to 45MB on Mip-NeRF360—a 15$\times$ reduction.
+LightGaussian also outperforms 3D-GS on the Tank & Temple datasets, nearly doubling rendering speed and reducing storage from 380MB to 22MB.
 Qualitative Results.
-We conducted a comparative analysis of the rendering outcomes between 3D-GS and our LightGaussian model, focusing particularly on the intricate details and background regions, as demonstrated in Fig.
+We conducted a comparative analysis of rendering results between 3D-GS and LightGaussian, focusing on intricate details and background regions, as shown in Fig.
 4.
-It is evident that both 3D-GS and LightGaussian maintain comparable visual quality, even in scenes with thin structures (refer to the first and third rows).
+Both 3D-GS and LightGaussian exhibit comparable visual quality, even in challenging scenes with thin structures, showing that LightGaussian effectively removes redundancy while preserving reconstruction fidelity.
+Generalization to Other Point-based Representations
+We further apply our Gaussian Pruning approach to Scaffold-GS [67], an advanced neural Gaussian-based 3D representation.
+In this method, we prune Gaussians that contribute minimally to scene reconstruction, reducing redundancy while preserving visual fidelity.
+Experiments on the MipNeRF360 dataset demonstrate that pruning 80% of neural Gaussians increases the rendering speed from 152 to 178 FPS, as shown in Tab.
+5.
+These results underscore LightGaussian’s potential as an effective optimization tool for other Gaussian-based representations.
+Size FPS PSNR SSIM LPIPS
+Scaffold-GS 173.60 152 27.96 0.8240 0.2075
+Scaffold-GS + LightGaussian 112.56 178 27.78 0.8187 0.2197
+Table 5: LightGaussian removes redundant neural Gaussians in Scaffold-GS, reducing model size and improving rendering speed.
+All experiments were rerun on our platform for fair comparison, with results averaged on the MipNeRF-360 dataset.
 4.3 Ablation Studies
-We conducted ablation studies on our method’s components to individually assess the impact of the proposed modules.
-Our findings reveal that Gaussian Pruning and Recovery is effective in eliminating redundant Gaussians, ensuring that only significant ones contribute to the scene representation.
-SH Distillation proves to be effective in reducing the degree of Spherical Harmonics, which simplifies the lighting model without compromising too much quality.
-Additionally, Vector Quantization demonstrates its efficiency in compressing the feature space, resulting in a more streamlined and efficient representation.
-When combined, these modules significantly enhance the overall efficacy and performance of our framework.
-Overall Analysis.
-We report the performance of the investigation on the proposed modules in Tab.
-2 and Fig.
+We performed ablation studies on each component of our method to evaluate their individual impacts.
+Results show that Gaussian Pruning and Recovery effectively removes redundant Gaussians, retaining only those crucial for scene representation.
+SH Distillation reduces the Spherical Harmonics degree, simplifying the lighting model with minimal quality loss.
+Furthermore, Vector Quantization efficiently compresses the feature space, creating a more compact representation.
+Combined, these modules substantially improve the overall efficiency and performance of our framework.
+Significance Criteria
+To identify the optimal criteria for measuring each Gaussian’s global significance, we evaluated key characteristics: Gaussian-ray interaction count, Gaussian opacity, and functional volume.
+Tab.
+3 illustrates that integrating these three factors — by weighting the hit count with opacity and Gaussian volume — yields the highest rendering quality post zero-shot pruning.
+A quick Gaussian Co-adaptation (the last row) which optimizes the remaining Gaussians, can effectively recover the rendering accuracy to its pre-pruning level.
+The visual comparison of rasterized images, before and after pruning, alongside the depiction of pruned Gaussians, is presented in Fig.
 5.
-We verify that our design of Gaussian Pruning & Recovery is effective in removing redundant Gaussians (Exp #1$\rightarrow$3) with negligible quality degradation while preserving rendering accuracy.
-This proves that the proposed global significance, based on principle of splatting, accurately represents the critical aspects of each Gaussian.
-By removing the high-degree SHs and transferring the knowledge to a compact representation (Exp #3 $\rightarrow$5), our method successfully demonstrates the benefits from using soft targets and extra data from view augmentation, results in negligible changes in specular reflection.
-In practical post-processing, the vector quantization on the least important Gaussians (Exp #7), showcases the advantage of adopting VQ to further reduce model size.
 Figure 5: Visualization of Pruned Gaussians.
@@
 The residual is visualized by rasterizing the pruned Gaussians.
-The Best Significance Criteria?
-To discover the most favorable criteria to measure the global significance of each Gaussian, we evaluated key characteristics within each Gaussian: the Gaussian/rays interaction count, Gaussian opacity, and the volume related to its function.
-Tab.
-3 illustrates that integrating these three factors—by weighting the hit count with opacity and Gaussian volume—yields the highest rendering quality post zero-shot pruning.
-A quick Gaussian Co-adaptation (the last row) which optimizes the remaining significant Gaussians, can effectively recover the rendering accuracy to its pre-pruning level.
-The visual comparison of rasterized images, before and after pruning, alongside the depiction of pruned Gaussians, is presented in Fig.
-6.
 SH Distillation & Vector Quantization.
-The direct removal of high-degree components in Spherical Harmonics (SHs) results in a substantial performance decline compared to the full model (Exp #3 $\rightarrow$ 4), manifesting notably in the loss of specular reflections on surfaces with varying viewpoints.
-However, the integration of knowledge distillation (Exp #4 $\rightarrow$ 5) from the full model enables us to decrease its size while maintaining the essential viewing effects.
-Furthermore, the incorporation of pseudo-views during the training process (Exp #5 $\rightarrow$ 6) validate the efficacy of using the simulated views to enhance the learning of specular reflections.
-Applying Vector Quantization (VQ) on all attributes leads to inferior results (See Tab.
-4), but this can be mitigated by using Gaussian Global Significance.
-Table 5: Ablation under different SH ompactness: With a Gaussian pruning ratio of 66% and reduce SH to 2-degree, LightGaussian nearly preserves the rendering quality (SSIM decrease from 0.927 to 0.926).
-Reducing the SH to 1-degree lead to the SSIM decreases to 0.923.
+Directly removing high-degree components in Spherical Harmonics (SHs) causes significant performance degradation compared to the full model (Exp #3 $\rightarrow$ 4), particularly with the loss of specular reflections across varying viewpoints.
+However, introducing knowledge distillation from the full model (Exp #4 $\rightarrow$ 5) allows for size reduction while preserving essential viewing effects.
+Additionally, incorporating pseudo-views during training (Exp #5 $\rightarrow$ 6) demonstrates the effectiveness of simulated views in enhancing the learning of specular reflections.
+While applying VQ to all attributes yields poorer results (see Tab.
+4), this impact is mitigated by leveraging Gaussian Global Significance.
 Exp ID Model setting Model Size $\downarrow$ FPS $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
@@
 [5] [4] + proposed distillation 88.89 MB 244 31.440 0.923 0.205
-Figure 7: Visualization of the rendering performance vs. different prune and VQ ratios. The quantize ratio is the compression ratio at vector quantization.
+Table 6: Ablation with varying SH compactness: With a Gaussian pruning ratio of 66% and SH reduced to 2 degrees, LightGaussian nearly maintains rendering quality (SSIM drops slightly from 0.927 to 0.926).
+Reducing SH further to 1 degree lowers SSIM to 0.923.
+Figure 6: Rendering performance comparison across different pruning levels and VQ ratios.
+Note that in subfigure (b), we use the VQ ratio to indicate the proportion of SH parameters quantized to an 8192-codebook, while the (overall) quantization ratio reflects combined compression from VQ and FP32-to-FP16 bitwise quantization.
 Degradation & Speed vs. Compression Ratio
@@
 In Tab 6, we observe that 2-degree SH compactness reduces the SSIM from 0.927 to 0.926, while 1-degree SH compactness further reduces it to 0.923.
-5 More Technical Details
+5 Conclusion, Limitations, and Broad Impact
+We present LightGaussian, a novel framework that transforms heavy point-based representations into a compact format for efficient novel view synthesis.
+Designed for practical use, LightGaussian leverages 3D Gaussians to model large-scale scenes, effectively identifying and pruning the least significant Gaussians generated through densification.
+It achieves over 15$\times$ data reduction, boosts FPS to over 200, and minimally impacts rendering quality.
+Exploring zero-shot compression across various 3D-GS-based frameworks remains a promising direction for future research.
+Broadly, LightGaussian’s compact representation has the potential to democratize high-quality 3D content for applications in VR, AR, and autonomous driving by reducing resource demands, enabling more accessible and scalable deployment across industries.
+While we do not foresee any significant risk from this specific technique, 3D reconstruction may infringe on personal privacy when applied in public spaces or with drone footage, contradicting our ethical intentions.
+Appendix
 Algorithm Explanation.
@@
 22:Save optimized model $\mathcal{G^{\prime\prime\prime}}$ to disk.
-Specifically, we describe how we calculate the Global Significance Score for each trained Gaussian in Algorithm 2.
+Specifically, we detail how we calculate each Gaussian’s Global Significance Score in Algorithm 2.
 Algorithm 2 Global significance calculation for gaussians.
@@
 This approach permits precise adjustments to the codebook and the rest attributes, improving its alignment with the training images (#8, SSIM improves from 0.915 to 0.923).
-6 Detailed Experiment Results
+6 More Experiment Results
 In addition to realistic indoor and outdoor scenes in the Mip-NeRF360 [7] and Tank and Temple datasets [65], we further evaluate our method on the synthetic Blender dataset [5], and provide a scene-wise evaluation on all datasets, accompanied by detailed visualizations.
@@
 Overall comparisons with previous methods are listed in Table 7, where we observe our LightGaussian markedly reduces the average storage size from 52.38MB to 7.89MB, while improving the FPS from 310 to 411 with only a slight rendering quality decrease.
+Figure 7: Visual Comparisons for Ablation Study.
+We visualize the rendered RGB images and the residual map between the ground-truth image, aligned with the experiment ID as shown in Tab.
+The final model (Exp #9) demonstrates close results to 3D-GS (Exp #1), while the Gaussian Co-adaptation, along with SH distillation, almost completely mitigates the information loss.
+To highlight the difference, we visualize the SSIM between the rendered images and the GT, where a whiter output in the residual signifies closer alignment with the GT, denoting better quality, and the area with greater color intensity indicates lower quality.
 Table 7: Per-scene results on Synthetic-NeRF.
@@
 PSNR(dB)
-3D-GS 35.436 26.294 35.614 37.848 35.782 30.533 36.585 31.642 33.716
+3D-GS 35.436 26.294 35.614 37.848 35.45 30.533 36.585 31.642 33.716
 Ours 34.769 26.022 34.484 36.461 34.944 29.341 35.370 30.405 32.725
@@
 Additional Quantitative Results on Mip-NeRF360.
-Tables 7, 8, and 9 present the comprehensive error metrics compiled for our evaluation across all real-world scenes (Mip-NeRF360 and Tank and Temple datasets).
-Our method not only compresses the average model size from 724MB to 42MB, but also consistently demonstrates comparable metrics with 3D-GS on all scenes.
+Tables 8, 9, and 10 present the comprehensive error metrics compiled for our evaluation across all real-world scenes (Mip-NeRF360 and Tank and Temple datasets).
+Our method not only compresses the average model size from 782MB to 45MB, but also consistently demonstrates comparable metrics with 3D-GS on all scenes.
 LightGaussian additionally shows better rendering quality than Plenoxel, INGP, mip-NeRF360, and VQ-DVGO.
@@
 We present the rendering results from 3D-GS [16], LightGaussian, and VQ-DVGO [49].
-The corresponding residual maps highlighting the differences between the rendered images and ground truth (GT) images are also displayed.
+The residual maps highlight the differences between the rendered images and ground truth (GT) images.
 Figure 9: Additional Visual Comparisons on the Mip-NeRF360 Datasets.
 We present the rendering results from 3D-GS [16], LightGaussian, and VQ-DVGO [49].
-The corresponding residual maps highlighting the differences between the rendered images and ground truth (GT) images are also displayed.
-Table 7: Quantitative Comparison (PSNR) on Mip-NeRF360 and Tank & Temple Scenes. We re-train 3D-GS using their original code and configurations on our platform, to perform a fair comparison.
+The residual maps highlight the differences between the rendered images and ground truth (GT) images.
+Table 8: Quantitative Comparison (PSNR) on Mip-NeRF360 and Tank & Temple Scenes.
+We re-train 3D-GS, Compressed 3D-GS, Compact 3D-GS using their original code and 3D-GS configurations on our platform, to perform a fair comparison.
 Methods PSNR
@@
 VQ-DVGO 22.092 18.934 24.127 23.428 28.502 21.440 26.035 25.459 27.990 - -
-3D-GS 25.122 21.324 27.294 26.783 31.687 22.500 29.114 31.618 32.307 24.778 21.449
-Ours 25.093 21.252 27.043 26.862 31.680 22.658 28.464 30.974 31.459 24.851 21.407
+3D-GS 25.194 21.414 27.29 26.598 31.336 22.476 28.989 31.328 31.948 25.372 21.957
+Compressed 3D-GS 25.040 21.126 26.817 26.357 31.072 22.298 28.653 30.755 31.161 25.194 21.882
+Compact 3D-GS 24.795 20.974 26.703 26.260 30.645 22.570 28.553 30.332 31.84 25.041 21.544
+Ours 25.196 21.538 26.961 26.77 31.399 22.685 28.478 30.869 31.414 25.399 21.838
+Table 9: Quantitative Comparison (SSIM) on Mip-NeRF360 and Tank & Temple Scenes.
+We re-train 3D-GS, Compressed 3D-GS, Compact 3D-GS using their original code and 3D-GS configurations on our platform, to perform a fair comparison.
+Methods SSIM
+bicycle flowers garden stump room treehill counter kitchen bonsai truck(T&T) train(T&T)
+Plenoxels 0.496 0.431 0.606 0.523 0.509 0.842 0.759 0.648 0.814 0.774 0.663
+INGP-Big 0.512 0.486 0.701 0.594 0.542 0.871 0.817 0.858 0.906 0.800 0.689
+mip-NeRF360 0.685 0.584 0.809 0.631 0.745 0.910 0.892 0.917 0.938 0.857 0.660
+VQ-DVGO 0.473 0.355 0.614 0.563 0.861 0.476 0.777 0.760 0.842 - -
+3D-GS 0.764 0.601 0.862 0.770 0.917 0.632 0.906 0.925 0.939 0.878 0.811
+Compressed 3D-GS 0.752 0.584 0.845 0.757 0.912 0.622 0.896 0.916 0.932 0.873 0.803
+Compact 3D-GS 0.737 0.568 0.837 0.753 0.909 0.632 0.892 0.913 0.931 0.869 0.788
+Ours 0.765 0.598 0.855 0.776 0.916 0.635 0.898 0.919 0.934 0.877 0.801
+Table 10: Quantitative Comparison (LPIPS) on Mip-NeRF360 and Tank & Temple Scenes.
+We re-train 3D-GS, Compressed 3D-GS, Compact 3D-GS using their original code and 3D-GS configurations on our platform, to perform a fair comparison.
+Methods LPIPS
+bicycle flowers garden stump room treehill counter kitchen bonsai truck(T&T) train(T&T)
+Plenoxels 0.506 0.521 0.3864 0.503 0.540 0.4186 0.441 0.447 0.398 0.335 0.422
+INGP-Big 0.446 0.441 0.257 0.421 0.450 0.261 0.306 0.195 0.205 0.249 0.360
+mip-NeRF360 0.305 0.346 0.171 0.265 0.347 0.213 0.207 0.128 0.179 0.159 0.354
+VQ-DVGO 0.571 0.628 0.402 0.425 0.216 0.640 0.244 0.222 0.191 - -
+3D-GS 0.211 0.339 0.108 0.216 0.221 0.327 0.201 0.127 0.206 0.147 0.209
+Compressed 3D-GS 0.237 0.358 0.139 0.248 0.234 0.352 0.215 0.140 0.219 0.157 0.220
+Compact 3D-GS 0.256 0.378 0.144 0.256 0.233 0.347 0.223 0.140 0.218 0.163 0.240
+Ours 0.218 0.352 0.122 0.222 0.232 0.338 0.220 0.141 0.221 0.155 0.239
 Implementation Details of VQ-DVGO.
@@
 All other parameters are aligned with those specified in the original VQ-DVGO paper [49], ensuring a faithful replication of established methodologies.
-Table 8: Quantitative Comparison (SSIM) on Mip-NeRF360 and Tank & Temple Scenes. We re-train 3D-GS using their original code and configurations on our platform, to perform a fair comparison.
-Methods SSIM
-bicycle flowers garden stump room treehill counter kitchen bonsai truck(T&T) train(T&T)
-Plenoxels 0.496 0.431 0.606 0.523 0.509 0.842 0.759 0.648 0.814 0.774 0.663
-INGP-Big 0.512 0.486 0.701 0.594 0.542 0.871 0.817 0.858 0.906 0.800 0.689
-mip-NeRF360 0.685 0.584 0.809 0.631 0.745 0.910 0.892 0.917 0.938 0.857 0.660
-VQ-DVGO 0.473 0.355 0.614 0.563 0.861 0.476 0.777 0.760 0.842 - -
-3D-GS 0.746 0.574 0.856 0.770 0.926 0.630 0.914 0.932 0.946 0.863 0.781
-Ours 0.745 0.566 0.845 0.775 0.923 0.627 0.900 0.921 0.938 0.862 0.772
-Table 9: Quantitative Comparison (LPIPS) on Mip-NeRF360 and Tank & Temple Scenes. We re-train 3D-GS using their original code and configurations on our platform, to perform a fair comparison.
-Methods LPIPS
-bicycle flowers garden stump room treehill counter kitchen bonsai truck(T&T) train(T&T)
-Plenoxels 0.506 0.521 0.3864 0.503 0.540 0.4186 0.441 0.447 0.398 0.335 0.422
-INGP-Big 0.446 0.441 0.257 0.421 0.450 0.261 0.306 0.195 0.205 0.249 0.360
-mip-NeRF360 0.305 0.346 0.171 0.265 0.347 0.213 0.207 0.128 0.179 0.159 0.354
-VQ-DVGO 0.571 0.628 0.402 0.425 0.216 0.640 0.244 0.222 0.191 - -
-3D-GS 0.245 0.381 0.122 0.242 0.200 0.366 0.186 0.118 0.184 0.175 0.262
-Ours 0.257 0.392 0.143 0.251 0.212 0.378 0.211 0.140 0.199 0.179 0.283
-7 Conclusion and Future Works
-We present LightGaussian, a novel framework that converts the heavy point-based representation into a compact format for efficient novel view synthesis.
-For practical use, LightGaussian explores the use of 3D Gaussians for modeling large-scale scenes and finds an effective way to identify the least important Gaussians grown by densification.
-In pursuing a compact Spherical Harmonics format, distillation is enhanced by synthesizing pseudo-views to generate more data for knowledge transfer.
-A 3D Gaussian Vector Quantization post-processing effectively removes further redundancy by using Gaussian significance as an indicator.
-With these methods, the proposed representation reduces data redundancy by more than 15$\times$, further boosting the FPS to more than 200 FPS, with minimal rendering quality loss.
-Exploring the possibility of zero-shot compression across various 3D-GS-based frameworks is a valuable direction for future research.
+Overall Analysis.
+We report the performance of the investigation on the proposed modules in Tab.
+2 and Fig.
+7.
+We verify that our design of Gaussian Pruning & Recovery is effective in removing redundant Gaussians (Exp #1$\rightarrow$3) with negligible quality degradation while preserving rendering accuracy.
+This proves that the proposed global significance, based on principle of splatting, accurately represents the critical aspects of each Gaussian.
+By removing the high-degree SHs and transferring the knowledge to a compact representation (Exp #3 $\rightarrow$5), our method successfully demonstrates the benefits from using soft targets and extra data from view augmentation, results in negligible changes in specular reflection.
+In practical post-processing, the vector quantization on the least important Gaussians (Exp #7), showcases the advantage of adopting VQ to further reduce model size.